### v2

# Think or Step-by-Step? UnZIPping the Black Box in Zero-Shot Prompts

Abstract
Zero-shot prompting techniques have significantly improved the performance of Large Language Models (LLMs). However, we lack a clear understanding of why zero-shot prompts are so effective. For example, in the prompt “Let’s think step-by-step,” is think or step-by-step more crucial to its success? Existing interpretability methods, such as gradient-based and attention-based approaches, are computationally intensive and restricted to open-source models. We introduce the ZIP score (Zero-shot Importance of Perturbation score), a versatile metric applicable to both open and closed-source models, based on systematic input word perturbations. Our experiments across four flagship models, seven widely-used prompts, and several tasks reveal that: (1) word importance varies significantly by task type, with mathematical tasks prioritizing structured instruction words; (2) proprietary models align more closely with human judgments than open-source ones; (3) nouns consistently show the highest importance (47.4%–65.9%); and (4) ZIP scores inversely correlate with model performance (|r| > 0.9), indicating stronger prompt effects on challenging tasks. These findings enhance our understanding of LLM behavior and contribute to developing more effective zero-shot prompts and improved model analysis. [^1] The complete codebase and documentation of language model interactions will be publicly available upon publication.
[…]
Think or Step-by-Step?
UnZIPping the Black Box in Zero-Shot Prompts
Nikta Gohari Sadr, Sangmitra Madhusudan, and Ali Emami Brock University, Saint Catharines, Canada {zu22of, sm20pd, aemami}@brocku.ca
1 Introduction
Figure 1: ZIP score generation process for the Chain-of-Thought prompt using GPT-4o mini on a AQUA-RAT dataset instance. Perturbed prompts are compared to the original, generating word-level ZIP scores. The red box highlights “step-by-step” as significantly important based on statistical analysis.
Prompting techniques have revolutionized how we use Large Language Models (LLMs), significantly improving their performance across various tasks. From Chain-of-Thought (CoT) Wei et al. (2022) to Self-Consistency Wang et al. (2023b) and Least-to-Most prompting Zhou et al. (2022), these methods have expanded LLM capabilities. Zero-shot prompts, which guide models without task-specific training examples, have emerged as particularly effective and practical, offering broad generalization and multi-task abilities Kojima et al. (2022). However, despite their success, we lack a clear understanding of why these prompts are so effective, particularly at the level of individual words.
Consider the following example:
Example 1.
Problem: Alice has 5 apples. She gives 2 to Bob and 1 to Charlie. How many apples does Alice have left?
Zero-shot CoT: Let’s think step-by-step.
Model Output:
  1. 1.
    Alice starts with 5 apples.
  2. 2.
    She gives away 2 + 1 = 3 apples.
  3. 3.
    Apples left = 5 - 3 = 2
Answer: Alice has 2 apples left.
Despite the success of such zero-shot prompting techniques, a gap in our understanding remains: What is it about these prompts that is so effective, particularly at the level of individual words? In the above prompt, which term has more influence on the model output: ‘think’ or ‘step-by-step’?
Current methods for interpreting input word significance in LLMs typically rely on gradients Wallace et al. (2019); Yin and Neubig (2022); Ferrando et al. (2023); Wu et al. (2023) or attention mechanisms Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022). However, these methods require access to internal model parameters and face reliability issues Wang et al. (2020); Jain and Wallace (2019); Ethayarajh and Jurafsky (2021); Zhao et al. (2024); Bastings and Filippova (2020).
Following established work Choudhary et al. (2022); Saleem et al. (2022); Feng et al. (2018), we define word importance as the degree to which a word’s presence, absence, or modification impacts model performance on a given task. This definition captures both semantic and structural importance through systematic perturbations: a word is considered important if replacing it with semantically similar alternatives (synonyms), related concepts (co-hyponyms), or removing it altogether significantly affects the model’s output.
Our paper introduces the ZIP score (Zero-shot Importance of Perturbation score), a novel metric for assessing word importance in zero-shot instructional prompts that works with both open and closed-source models. Figure 1 provides an overview of the methodology. Our approach is validated through controlled experiments, statistical analyses, and is compared with human judgment across seven zero-shot prompts, four LLMs, and several classification and translation tasks.
Our key findings reveal that: (1) Word importance varies significantly by task type (e.g., "step-by-step" for mathematical tasks vs. "think" for reasoning tasks); (2) Proprietary models like GPT-4o mini align more closely with human judgments of word importance than open-source models; (3) Nouns consistently show the highest importance (47.4%–65.9% of significant words) across all models; and (4) ZIP scores inversely correlate with model performance (|r| > 0.9), indicating that prompts have a greater impact on challenging tasks.
2 Related Work
Feature importance interpretation in language models is typically approached through three main methods: gradient-based, attention-based, and perturbation-based Choudhary et al. (2022).
Gradient-Based Methods identify influential features by calculating the gradient of the output logits with respect to input elements Wallace et al. (2019). Recent advances include gradient-based post hoc contrastive explanations for model predictions Yin and Neubig (2022); Ferrando et al. (2023) and analyses of how CoT prompting affects saliency scores Wu et al. (2023). However, these methods face significant limitations: they require access to model internals, making them inapplicable to closed-source models. Furthermore, research has shown that model gradients can be easily manipulated, raising concerns about the reliability of gradient-based analyzes Wang et al. (2020).
Attention-Based Methods interpret feature importance by analyzing the weighted sum of intermediate representations in neural networks Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022). While intuitive, these approaches have several drawbacks. For one, the attention weights do not always correspond to the most important features for model predictions Jain and Wallace (2019) and may not correlate well with other feature importance measures Ethayarajh and Jurafsky (2021). Moreover, attention weights often contain redundant and mixed information, leading to unreliable explanations Bastings and Filippova (2020). Like gradient-based methods, these approaches also require access to model internals, limiting their applicability to open-source models.
Perturbation-Based Methods can analyze any LLM regardless of architecture accessibility by measuring how outputs change when inputs are modified. As Choudhary et al. (2022) describes, “a word (token) or a collection of words (tokens) are modified or removed from the input samples, and a resulting change is measured.”
Notable approaches include LIME Ribeiro et al. (2016), which creates local linear approximations of model behavior, and other dataset instance perturbations Zafar and Khan (2019). Recent research has extended these methods to few-shot demonstrations Liu et al. (2023) and system prompts Hackmann et al. (2024); Yin et al. (2023).
Despite these advancements, there remains a significant gap in the interpretation of zero-shot instructional prompts. Current research often uses basic token masking, which may oversimplify prompt semantics. It can also result in incoherent perturbations, potentially failing to interpret the true impact of the word on model predictions Yin et al. (2023); Feng et al. (2018). Our research addresses this by introducing a new metric that employs a variety of meaningful perturbations to reveal the importance of each word in zero-shot instructional prompts.
3 The ZIP score
The ZIP score quantifies the importance of individual words in a prompt by measuring how much the model’s performance changes when that word is perturbed or removed. This method allows us to systematically identify which words are most crucial for the model’s performance on a given task.
3.1 Formalization of the ZIP score
We formalize the ZIP score through the following steps, using a running example:
Example Task ($T$): Determine whether a number is prime or composite.
Original Prompt ($P$): “Take a deep breath and work on this problem step-by-step” Luo et al. (2024)
Step 1: Prompt Representation
Let $P$ represent the original prompt, where $P=w_{1},w_{2},\dots,w_{n}$ and $w_{i}$ denotes the $i$-th word. [^2] Throughout this paper, we use “word” to refer to any token resulting from our space-based tokenization method. This may include contractions, compound words, or multi-word expressions.
Example:
[…]
Figure 2: Perturbation generation and filtering process for a sample prompt
Step 2: Perturbation Generation
For each word $w_{i}$ in $P$, we generate a set of perturbations $\{P_{i1},P_{i2},\dots,P_{ik}\}$, through context-aware synonym replacement and co-hyponym replacement using GPT-4333GPT-4 consistently outperformed traditional tools like NLTK Bird et al. (2009) in generating contextually appropriate alternatives in our preliminary tests., and word removal. Synonyms provide similar meanings, co-hyponyms [^4] A co-hyponym is a word that shares the same category or group with another word, but both are specific examples of a broader term (hypernym). introduce related concepts, and removal tests word necessity. Each $P_{ij}$ represents a modified version of $P$ where $w_{i}$ has been replaced or omitted.
[…]
Aligned with existing literature on perturbation-based methods Choudhary et al. (2022); Ivanovs et al. (2021); Saleem et al. (2022), we posit that a word’s importance is indicated by the model’s performance shifts in response to any meaningful alteration of that word, whether through subtle meaningful shifts or its complete removal. (see §7 for analyses and demonstrations of this phenomenon).
To create a final set of perturbations that provides meaningful insight into the impact of each word on the model output, as depicted in Figure 2, we filter the perturbed sentences as follows:
  1. 1.
    The Universal Sentence Encoder Cer et al. (2018) to measure semantic similarity between $P$ and $P_{ij}$, keeping those with >30% similarity. [^5] This threshold was chosen to balance maintaining context while allowing meaningful alterations (see table 24 in A.7).
  2. 2.
    GPT-4 evaluation to assess grammatical correctness and semantic viability of each $P_{ij}$.
Determined by our filtering process, $P_{42}$ is excluded due to its insufficient semantic similarity and grammatical inconsistencies. Consequently, the final set of perturbations includes $P_{41}$ and $P_{43}$.
We maintained cost-effectiveness in using large language models by optimizing token usage and the output format, reducing the tokens needed for perturbation generation and selection. Additionally, our primary filter, using Universal Sentence Encoder, discards unrelated perturbations, thereby minimizing unnecessary API calls for grammatical verification [^6] For the CoT prompt, the entire perturbation generation and grammatical verification process uses fewer than 2150 input/output tokens, effectively balancing cost and quality.. Across the seven instructional prompts, this filtering process resulted in an average of 9.04 perturbed sentences per word. Detailed prompts and examples for perturbation creation and evaluation are provided in Appendix Sections A.6 and A.7.
Step 3: Model Prediction
We input the original prompt $P$ and each of its validated perturbed variants $P_{ij}$ into the language model $M$ to address an instance of task $T$ which we refer to as $t$.
Example ($t$): Is 29 prime or composite?
[…]
Step 4: Disagreement Calculation
We define a disagreement function $d$ that measures the difference between the model predictions for each perturbed prompt ($P_{ij}$) and the original prompt ($P$). This function varies according to the type of task $T$. For instance $t$ of classification tasks:
$d_{t}^{M}(P,P_{ij})=\begin{cases}1&\text{if }\text{pred}_{t}^{M}(P)\neq\text{pred}_{t}^{M}(P_{ij})\\ 0&\text{otherwise}\end{cases}$
Example:
[…]
For an instance $t$ of translation tasks:
$d_{t}^{M}(P,P_{ij})=|\text{BLEU}_{t}^{M}(P)-\text{BLEU}_{t}^{M}(P_{ij})|$
Here $\text{BLEU}_{t}^{M}(P)$ and $\text{BLEU}_{t}^{M}(P_{ij})$ are the BLEU scores Papineni et al. (2002) for translations generated from the original and perturbed prompts, respectively.
Translation Example: For the task of translating “The cat is on the mat” to French, assume:
$\text{BLEU}_{t}^{M}(P)=0.85$,
$\text{BLEU}_{t}^{M}(P_{41})=0.83$,
$\text{BLEU}_{t}^{M}(P_{43})=0.72$
$d_{t}^{M}(P,P_{41})=|0.85-0.83|=0.02$, $d_{t}^{M}(P,P_{43})=|0.85-0.72|=0.13$
Step 5: ZIP score Calculation Finally, for each word $w_{i}$, we compute the ZIP score $\text{ZIP}_{T}^{M}(w_{i})$ by averaging disagreement scores across all perturbations of $w_{i}$ and all dataset instances:
$\text{ZIP}_{T}^{M}(w_{i})=100\cdot\frac{1}{N}\sum_{t=1}^{N}\left(\frac{1}{k}\sum_{j=1}^{k}d_{t}^{M}(P_{i},P_{ij})\right)$
Where $N$ is the number of dataset instances, $k$ is the number of valid perturbations for each word, $P_{i}$ is the original prompt for the $i$-th dataset instance, and $P_{ij}$ is the $j$-th perturbation of word $w_{i}$ in the $t$-th dataset instance.
[…]
$\text{ZIP}_{Classif.}^{M}(\text{``breath''})=100\cdot\frac{1}{2}(0+1)=50$
$\text{ZIP}_{Transl.}^{M}(\text{``breath''})=100\cdot\frac{1}{2}(0.02+0.13)=7.5$
The ZIP score ($\in[0,100]$) [^7] The 100x scaling factor normalizes ZIP scores to a 0-100 range for easier interpretation and comparison across tasks. represents the average percentage change in model performance observed across all tested dataset instances when perturbing a given word. [^8] For simplicity, we show the process for one word in a dataset instance; in practice, this repeats for all prompt words, averaging ZIP scores across instances. This metric is adaptable to various performance measures, making it versatile for different types of tasks. Interpreting the ZIP score is straightforward: a higher score indicates greater word importance, as perturbing it leads to larger changes in model output. Conversely, a lower score suggests the word has less impact on performance for the given task.
3.2 Identifying Significantly Important Words
To determine which words are significantly important in prompts, we use a statistical approach that distinguishes between the inherent variability in model outputs and the effects of our perturbations:
  1. 1.
    For each word in the prompt, we generate a set of perturbed prompts (the same set used in the ZIP score calculation).
  2. 2.
    We repeatedly input both the original prompt and each of its perturbed versions to the model, generating multiple outputs for each. The number of times we re-prompt with the original prompt matches the number of perturbations for that word, ensuring a balanced comparison.
  3. 3.
    We compare the ZIP scores from the original re-prompts with those from each perturbed prompt using the Wilcoxon rank-sum test.
The Wilcoxon rank-sum test evaluates whether two independent samples (i.e., the outputs from the original re-prompts and a perturbed prompt) come from the same distribution. Words yielding a p-value < 0.05 are considered ‘significantly important’.
4 Controlled Pre-Experiment
Despite the success of zero-shot prompts, establishing a ground truth for word-level importance within them is challenging because we may not know which words should matter most. To address this, we created a set of 20 simple zero-shot instructional validation prompts, each designed so that a single predetermined ‘key word’ determines the correct output. For example, consider the prompt “Say the word green”. Here, “green” is the key word: if the model does not produce “green” exactly, it fails. By knowing in advance which word is crucial for a correct output, we can objectively assess the effectiveness of interpretability methods.
Prompt Design and Methodologies: We constructed 20 validation prompts (Table 2) with varied syntax and structure, such as “Print the digits 123, “Type the letter X, and “Print carrot with no additional text”. Each prompt contains a key word that directly determines the correct output - any alteration to this word should change the model’s response. For example, in “Print the digits 123”, only perturbations to “123” should affect whether the model outputs “123” correctly. Table 11 (A.2) confirms these prompts isolate the key word’s influence on GPT-4o mini’s output.
We evaluate both our ZIP method and LIME Ribeiro et al. (2016), a widely-used perturbation-based explanation method, on this validation set. For ZIP, we apply our standard perturbation techniques to each validation prompt, measuring output changes using our disagreement function ($d=1$ for incorrect outputs, $d=0$ for correct ones). Unlike ZIP’s diverse perturbation approach, LIME relies solely on random word removal to estimate importance through linear modeling. Implementation details for LIME are provided in §5.
Validation Criteria: We consider an interpretability method successful if it consistently identifies the known key word as the most important term. For each prompt, a method scores correctly if it assigns the highest importance to the key word, and its accuracy is calculated as the percentage of prompts in which it succeeds. This provides a quantitative measure of the reliability of each method in identifying crucial prompt words.
5 Experimental Setup
Prompts
We evaluated seven zero-shot instructional prompts across various datasets, as detailed in Table 1. These prompts were selected based on their proven effectiveness and generalization capabilities in recent literature. We used a simple space-based tokenization method, where tokens are defined as sequences of characters separated by whitespace. [^9] This approach preserves contractions and compound words (e.g., “Let’s” as one token), which are often semantically important in instructional prompts. Dataset-specific prompts are included in the Appendix A.6 (Table 21).
Code Prompt Dataset
0-CoT Let’s think step-by-step. Kojima et al. (2024) MultiArith, GSM8K, AQUA-RAT, SVAMP, BIG-bench, etc.
0-CoTB Take a deep breath and work on this problem step-by-step. Yang et al. (2024) MultiArith, GSM8K, AQUA-RAT, BBH
0-CoTR Let’s work this out in a step-by-step way to be sure we have the right answer. Zhou et al. (2023) MultiArith, GSM8K, BIG-bench, etc.
0-IRR Feel free to ignore irrelevant information in the problem description. Shi et al. (2023) GSM-IC, GSM8k
0-PS Let’s first understand the problem and devise a plan to solve it. Then, solve it step-by-step.
(“ Plan & Solve”) Wang et al. (2023a)
MultiArith, GSM8K, AQUA-RAT, AddSub, SingleEq, etc.
0-DSP Provide the translation step-by-step, then complete the sentence. (“Domain Specific Prompting”) Peng et al. (2023) WMT22, WMT19 (German and Chinese)
0-DTG Detect the error type first, then refine the translation. (“Deliberate then Generate” Li et al. (2023) WMT, etc. (German, Czech, Chinese, etc)
Table 1: Overview of the zero-shot instructional prompts used in our experiments, alongside the datasets they were tested on. Datasets highlighted in bold indicate specific use for our study.
Datasets
We selected a diverse range of datasets known for effectively testing zero-shot prompts, as highlighted in bold in Table 1. For classification tasks, datasets included GSM8K Cobbe et al. (2021), AQUA-RAT Ling et al. (2017), and the Big Bench dataset Srivastava et al. (2023). Translation tasks used the WMT19 dataset, concentrating on Chinese and German languages. We evaluated on 150 instances per dataset, which we determined was sufficient to provide statistically significant results while managing computational costs.
Models
Four models were used in our experiments: GPT-4o mini, GPT-3.5-turbo, Llama-2-70B-chat, and Mixtral-8x7B-Instruct-v0.1 OpenAI (2023); Touvron et al. (2023); Jiang et al. (2024). Each model was configured with a temperature of 0.5 to balance between deterministic outputs and allowing for some variability. For token usage estimates per prompt, see Appendix A.8.
LIME Baseline
For our validation prompts, we configured LIME to treat the task as binary classification: Label 1 represents the target word (e.g., “green” for “Say the word green”), and Label 2 represents any other output. LIME automatically creates perturbations and fits a linear surrogate model to quantify word importance, using 50 perturbations per instance. This implementation allows direct comparison with ZIP scores while maintaining LIME’s standard approach to interpretability.
Human Evaluation
We conducted a study with 15 diverse participants proficient in English to compare human and model interpretations of the instructional language. For each prompt, participants selected up to three words that they deemed most important to solve a certain task. The evaluation form is shown in Appendix A.4 Fig. 17. We used the Jaccard Index to measure interannotator agreement across participants, yielding an average of 0.4714, indicating moderate agreement. [^10] Jaccard Index was chosen over Fleiss’ Kappa to accommodate multiple word selections per participant.
6 Results
6.1 Does the ZIP score accurately identify known important words?
Table 2 compares how ZIP and LIME performed in identifying key words on our 20 validation prompts on GPT-4o mini. ZIP correctly identified the predetermined key word in 90% of prompts, significantly outperforming LIME’s 60% accuracy. LIME’s errors often involved misidentifying contextual words (e.g., “Print,” “Output,” “Say") as most important, while ZIP maintained focus on the target words that directly determine the output.
ZIP’s strong performance was consistent across all four models (e.g., with 100% accuracy on GPT-3.5-turbo), validating ZIP’s effectiveness in identifying significant words in instructional prompts. Detailed results are provided in Appendix A.2.
Validation Prompts ZIP LIME
Say the word green Green Green
Print the digits 123 123 Print
Output the color blue Blue Output
Print carrot with no additional text Print Carrot
Display the word circle Circle Circle
When you’re ready just say coffee Coffee Coffee
Return the value five Five Five
Respond with only the word hello Hello Hello
Repeat the term mirror Mirror Repeat
Nine is the number you should write Nine Nine
Answer with the word pizza Pizza With
Begin by writing hello and then finish Hello Hello
Type purple in your response now Purple Purple
Say red and then stop Red Say
When you respond lead with river River River
Write the number seven Seven Number
Carefully type silver when responding here Silver Silver
Enter the word tomato Tomato The
Can you mention the direction up in our chat? Direction Chat
Type the letter X X X
Accuracy 90% 60%
Table 2: Most important words as identified by ZIP and LIME for GPT-4o mini on validation prompts. Red indicates failure to identify the correct key word.
6.2 How do ZIP scores vary across models for the CoT prompt?
Table 3 presents ZIP score heatmaps for the CoT prompt “Let’s Think Step-by-Step” across models on Big Bench. Our analysis reveals:
  • “Step-by-Step” vs “Think”: While both words are important, step-by-step shows, on average, a higher ZIP score across models. This suggests that the explicit instruction for a structured approach (“step-by-step”) may be more universally interpreted and utilized by models than the more general cognitive instruction (“think”).
  • Proprietary vs Open-source models: Proprietary models (GPT-4o mini, GPT-3.5-turbo) identify fewer words as significantly important (1 word) compared to open-source models (Mixtral, Llama 2, with 2-3 words). This suggests that proprietary models may be more selective in sensitivity to word perturbations, while open-source models appear to be sensitive to a wider range.
Results on all datasets are provided in Appendix A.1 Table 5.
GPT-4o mini [Uncaptioned image]
GPT 3.5 Turbo [Uncaptioned image]
Mixtral [Uncaptioned image]
Llama 2 [Uncaptioned image]
Table 3: ZIP score heatmaps for the CoT prompt across four models on the Big Bench dataset. Red boxes indicate significantly important words.
6.3 How does word importance vary across different prompts and task types?
AQUA-RAT Big Bench GSM8K
Top 3 MSWs ZIP Top 3 MSWs ZIP Top 3 MSWs ZIP
0-CoT Step-by-step 28.44 Think 3.80 Step-by-step 5.83
0-CoTB Step-by-step 34.49 Problem 1.95 Step-by-step 6.57
Work 1.33
0-CoTR Answer 27.66 Right 4.66 Answer 5.33
Sure 27.61 Work 4.40 Step-by-step 5.22
Right 26.44 Sure 4.09 Way 5.21
0-IRR Description 30.71 Ignore 3.87 Irrelevant 6.51
Ignore 30.42 Description 3.74 Description 6.41
Irrelevant 30.30
0-PS Plan 28.50 Plan 4.05 Step-by-step 7.28
Step-by-step 28.40 First 3.94 Problem 6.83
Solve 27.72 Solve 3.72 Solve 6.54
WMT 19: German WMT 19: Chinese
Top 3 MSWs ZIP Top 3 MSWs ZIP
0-DSP Translation 10.92 Translation 6.57
Step-by-step 7.68 Step-by-step 5.50
Sentence 6.50 Sentence 5.35
0-DTG Refine 8.43 Refine 5.31
Error 8.33 Type 4.92
Detect 8.06 Error 4.92
Table 4: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on GPT-4o Mini, with the most significant word in bold. All reported words are confirmed as significantly important.
Table 4 presents the ZIP score results for various zero-shot instructional prompts (detailed in Table 1) using the GPT-4o mini model. Our analysis demonstrates distinct patterns in the importance of words across various datasets and task types:
  • For mathematical and algebraic tasks (AQUA-RAT and GSM8K), the phrase step-by-step consistently emerges as highly important.
  • In contrast, for common sense reasoning tasks (Big Bench), words like think and problem show higher importance.
  • For translation tasks, translation and refine are identified as the most important words, which demonstrates the significance of task-specific instructions in multilingual contexts.
  • ZIP scores strongly correlate inversely with model performance (|r| > 0.9) across all prompts. For example, using CoT, GPT-4o mini shows higher ZIP scores on AQUA-RAT (accuracy: 68.26%, ZIP: 28.44) compared to GSM8K (91.59%, 5.83) and Big Bench (96.93%, 3.80) indicating that instructional prompts have greater impact on challenging tasks (Table 15).
Results for other models are provided in Appendix A.3 Tables 12 - 14.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 4: Part-of-Speech (POS) distribution of significantly important words across four LLMs.
6.4 How does the part-of-speech distribution of important words vary across models?
Figure 4 illustrates the part-of-speech (POS) distribution of the most important words for all seven instructional prompts. Key observations are:
  • Noun dominance: Nouns (NN) consistently rank as most important across all models (47.4%–65.9% of significant words).
  • Model patterns: GPT-4o mini and GPT-3.5-turbo, sharing similar foundational architectures despite vastly different parameter sizes, show similar POS distributions; Mixtral and Llama 2 show more balanced POS importance.
  • Verb importance: Base form verbs (VB) consistently rank 2nd in importance across all models.
  • Inter-model variation: Adverbs (RB) and non-3rd person singular present verbs (VBP) show the highest variation in importance across models.
6.5 How well do model assessments of word importance align with human judgments?
Figure 5 presents radar plots comparing human judgments of word importance with model-derived ones for the CoT prompt. Human assessment values represent the proportion of times each word was selected as important across all evaluators and tasks, while those of models indicate the proportion of times each word was determined as significantly important across all tasks.
Proprietary models (GPT-4o mini and GPT-3.5-turbo) align strongly with human judgments, while open-source models show less alignment, particularly Llama 2. “Step-by-step” emerges as highly important across all evaluations, while “Let’s” is consistently less emphasized. Similar analyses for remaining prompts are provided in Appendix A.4 Figures 11 - 16.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 5: Comparison of human judgments vs model-derived word importance for the CoT prompt.
7 Qualitative Analysis
We examined how subtle changes alter GPT-4o mini’s responses across tasks. In classification tasks, replacing “irrelevant” with “unrelated” in “Feel free to ignore irrelevant information” caused the model to overlook critical textual information, including numbers expressed in word form, leading to computational errors. For “Take a deep breath and work on this problem step-by-step", removing “step-by-step” surprisingly improved the model’s accuracy. In translation tasks, replacing “refine” with “enhance” in “Please detect the error type, and refine the translation” degraded translation quality, showing how synonym substitution can significantly affect task performance.
These examples demonstrate that seemingly equivalent words can produce significantly different model outputs. This observation supports our dual approach of using both synonym/co-hyponym substitution as well as word removal to reveal different aspects of how models interpret prompt wording. Details and heatmaps are provided in Appendix A.5 Tables 16 - 19.
8 Conclusion
This study introduces the ZIP score, a novel metric for quantifying word importance in zero-shot instructional prompts. Our findings reveal interesting patterns in word importance across models and tasks. For instance, while both ‘step-by-step’ and ‘think’ show high ZIP scores, which one is more influential depends on the model and task. We observed significant differences in word importance and alignment with human intuition between proprietary and open-source models. These insights aid our understanding of LLM behavior and open new avenues for prompt engineering and model analysis. Future work could explore the use of these findings to design more effective prompts tailored to specific model architectures.
Limitations
Scope of Perturbation Types: Our study strategically employed perturbations involving removals and semantic changes such as synonyms and co-hyponyms to analyze their impact on LLMs. While this approach provided valuable insights, extending the variety of perturbation types could uncover further nuances in language model behavior.
Interdependencies Among Prompt Constituents: This study evaluates the influence of individual words within prompts through isolation-based, a necessary simplification grounded in established practices to assess the independent significance of each word. However, it does not account for the interdependencies and contextual relationships between words that could collectively influence the output of LLMs. Future research should consider methods that can analyze the joint effects of multiple words, providing deeper insights into the model’s processing mechanisms, and improving the reliability of perturbation-based interpretability methods.
Model and Task Selection: This study concentrated on four models with different architectures selected for their capabilities and affordability. Although these models represent a robust cross section of current technologies, expanding this selection could improve the generalizability of our findings. Moreover, while our methods were applied to classification and translation tasks, they have potential for broader applications across different LLM tasks such as summarization or question answering, with only minor adjustments needed for the disagreement function to accommodate different outputs.
Sample Size Constraints: Our analysis, conducted on a carefully chosen sample of 150 instances from each dataset, was designed to balance depth with computational feasibility. Although this approach yielded statistically significant insights, increasing the sample size could further enhance the robustness and general applicability of our findings, capturing more subtle but potentially revealing perturbation effects.
Focus on Enhancing Interpretability: The primary aim of our research is not to engineer or optimize prompts to enhance the performance of LLMs, but rather to improve the interpretability of LLMs by analyzing existing effective prompting techniques. In future work, there is a valuable opportunity to leverage the insights gained from our interoperability method to craft new prompts explicitly designed to improve model performance.
Ethical Considerations in Application of Interpretability Insights: Although our method offers a deeper understanding of LLM operations, it is important to consider the potential for misuse. Detailed knowledge of how LLMs weigh different prompt elements could potentially be used to tailor content in ways that could manipulate or bias decision-making processes. It is crucial to apply these insights ethically with a commitment to fairness.
Appendix A Appendix
A.1 Chain-of-Thought ZIP score Heatmaps
GPT-4o mini Refer to caption
GPT 3.5 Turbo Refer to caption
Mixtral Refer to caption
Llama 2 Refer to caption
GPT-4o mini Refer to caption
GPT 3.5 Turbo Refer to caption
Mixtral Refer to caption
Llama 2 Refer to caption
Table 5: ZIP score heatmaps of the Chain-of-Thought (CoT) prompt across four models using the AQUA-RAT and GSM8K datasets. The tables display the significant word identification performance for each model with different datasets. Red boxes highlight words identified as significantly important, confirmed through statistical analysis.
A.2 Controlled Pre-Experiment
Validation Prompts ZIP LIME
Say the word green. Green Green
Print the digits 123. 123 123
Output the color blue. Blue Color
Print carrot with no additional text. Print/Carrot Carrot
Display the word circle. Cricle Word
When you’re ready just say coffee. Coffee Say
Return the value five. Five Value
Respond with only the word hello. Hello The
Repeat the term mirror. Mirror Mirror/Repeat
Nine is the number you should write. Nine Nine
Answer with the word pizza. Word/Pizza With
Begin by writing hello and then finish. Hello Hello
Type purple in your response now. Purple Purple
Say red and then stop. Red Red
When you respond lead with river. River Lead
Write the number seven. Seven Number
Carefully type silver when responding here. Silver Silver
Enter the word tomato. Word/Tomato The
Can you mention the direction up in our chat? Up Up
Type the letter X. X X
Accuracy 100% 55%
Table 6: Most important words as identified by ZIP and LIME for GPT 3.5 Turbo on validation prompts. Red indicates failure to identify the correct key word.
Validation Prompts ZIP LIME
Say the word green. Green Green
Print the digits 123. 123 Digits
Output the color blue. Color/Blue Blue
Print carrot with no additional text. Carrot Carrot
Display the word circle. Circle Circle
When you’re ready just say coffee. Coffee Coffee
Return the value five. Five Five
Respond with only the word Hello. Hello Hello
Repeat the term mirror. Mirror Term
Nine is the number you should write. Nine Nine
Answer with the word pizza. Pizza Pizza
Begin by writing hello and then finish. Hello Hello
Type purple in your response now. Purple Purple
Say red and then stop. Red Red
When you respond lead with river. River Respond
Write the number seven. Seven Seven
Carefully type silver when responding here. Silver Silver
Enter the word tomato. Tomato Tomato
Can you mention the direction up in our chat? Direction Chat
Type the letter X. X X
Accuracy 95% 80%
Table 7: Most important words as identified by ZIP and LIME for Mixtral on validation prompts. Red indicates failure to identify the correct key word.
Validation Prompts ZIP LIME
Say the word green. Green Green
Print the digits 123. Print/Digits/123 Digits
Output the color blue. Blue Blue
Print carrot with no additional text. Carrot No
Display the word circle. Circle Display
When you’re ready just say coffee. Coffee Say
Return the value five. Five Value
Respond with only the word Hello. Hello Hello
Repeat the term mirror. Mirror Mirror
Nine is the number you should write. Nine Should
Answer with the word pizza. Pizza Pizza
Begin by writing hello and then finish. Writing Hello
Type purple in your response now. Purple Purple
Say red and then stop. Red Red
When you respond lead with river. River River
Write the number seven. Seven Number
Carefully type silver when responding here. Silver Here
Enter the word tomato. Tomato Tomato
Can you mention the direction up in our chat? Chat Direction
Type the letter X. X X
Accuracy 90% 55%
Table 8: Most important words as identified by ZIP and LIME for Llama 2 on validation prompts. Red indicates failure to identify the correct key word.
GPT-4o mini [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GPT 3.5 Turbo [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Llama 2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Mixtral [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 9: Heatmap visualization of LIME for five validation prompts on GPT-4o mini, GPT 3.5 Turbo, Llama 2, and Mixtral models. The color blue represents alignment with Label 1 (the desired target word), and orange indicates Label 2 (the second most probable alternative output), with the intensity of each color indicating the level of importance attributed by LIME. Check (✓) and cross (✗) marks indicate whether the model correctly identified the predefined most important word.
GPT-4o mini [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GPT 3.5 Turbo [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Llama 2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Mixtral [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 10: Heatmap visualization of ZIP method results for five validation prompts, applied to GPT-4o mini, GPT 3.5 Turbo, Llama 2, and Mixtral models. These heatmaps illustrate the method’s ability to identify the most significant words across different prompts. Red boxes indicate significantly important words. Check (✓) and cross (✗) marks indicate whether the ZIP method correctly identified the predefined most important word.
Prompt Identified MSW Model Output
Print the digits 123. 123 123
Display the digits 123. 123
Print the digits _. 1234567890
Print the numbers 123. 123
Say the word green. Green Green
Say the word blue. Blue
Say the term green. Green
Say the word _. Hello
Repeat the term mirror. Mirror Mirror
Recite the term mirror. Mirror
Repeat the term glass. Glass
Repeat the word mirror. Mirror
Display the word circle. Circle Circle
Display the letter circle. O
Display the word Square. square
Display the word Round. round
Prompt Identified MSW Model Output
Print the digits 123 Print 123
_ the digits 123. One
_ _ digits _. Numbers
Print _ _ 123 123
Say the word green. Green Green
_ _ _ green Echo
Say the _ green Green
Say _ _ green Green
Repeat the term mirror. Repeat Mirror
_ the term mirror. Reflection
_ _ _ mirror. Reflect
Repeat _ term mirror. Mirror
Display the word circle. Circle Circle
_ _ word circle. Round
Display the word _. Do
Display _ word circle. Circle
Table 11: Comparison of GPT-4o Mini’s responses to the original validation prompt (bolded) and perturbed examples (underlined alterations), using ZIP and LIME methods. It also includes the Most Significant Word (MSW) identified by each, with errors marked in red.
A.3 ZIP Scores
AQUA-RAT Big Bench GSM8K
Top 3 MSWs ZIP Top 3 MSWs ZIP Top 3 MSWs ZIP
0-CoT Step-by-step 43.94 Step-by-step 48.38 Step-by-step 18.50
0-CoTB Problem 38.09 Step-by-step 38.71 Step-by-step 16.23
Breath 37.52 Breath 38.57
Problem 38.33
0-CoTR Sure 41.04 Right 39.59 Step-by-step 16.16
Sure 39.52
Let’s 38.88
0-IRR Feel 56.00 Description 49.43 Irrelevant 39.74
Description 50.15 Information 48.51 Ignore 38.06
Free 50.00 Free 48.22 Problem 35.09
0-PS Let’s 38.72 Plan 45.17 Devise 19.73
Carry 38.33 Problem 45.16 Plan 19.38
Problem 38.33 Let’s 45.08
WMT 19: German WMT 19: Chinese
Top 3 MSWs ZIP Top 3 MSWs ZIP
0-DSP Step-by-step 11.24 Step-by-step 7.90
Translation 8.78 Translation 6.92
Provide 7.39 Sentence 6.48
0-DTG Firstly 10.78 Firstly 7.34
Detect 10.04 Error 6.46
Translation 9.59 Detect 6.10
Table 12: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on GPT 3.5 Turbo, with the most significant word in bold. All reported words are confirmed as significantly important.
AQUA-RAT Big Bench GSM8K
Top 3 MSWs ZIP Top 3 MSWs ZIP Top 3 MSWs ZIP
0-CoT Let’s 58.53 Let’s 58.60 - -
Step-by-step 54.94 Step-by-step 58.50
0-CoTB Breath 58.38 Step-by-step 85.09 Step-by-step 33.66
Problem 57.47 Work 57.11 Breath 33.23
Take 57.11 Breath 54.66 Work 32.66
0-CoTR Step-by-step 55.22 Let’s 57.24 Step-by-step 32.16
Let’s 53.73 Work 57.00 Let’s 31.95
Way 50.42 Step-by-step 55.38 Work 29.93
0-IRR Ignore 46.48 Ignore 24.66 Description 45.12
Free 45.55 Irrelevant 21.23 Ignore 43.93
Description 42.92 Information 20.76 Irrelevant 43.02
0-PS Let’s 58.63 Carry 60.66 Understand 40.23
Solve 56.87 Understand 60.38 First 39.83
First 56.61 First 60.27 Problem 39.22
WMT 19: German WMT 19: Chinese
Top 3 MSWs ZIP Top 3 MSWs ZIP
0-DSP Translation 13.39 Step-by-step 7.55
Step-by-step 12.25 Translation 2.73
Following 12.02 Provide 2.34
0-DTG Detect 13.59 Please 8.00
Please 13.28 Type 7.91
Firstly 13.01 Firstly 7.89
Table 13: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on Mixtral, with the most significant word in bold. All reported words are confirmed as significantly important.
AQUA-RAT Big Bench GSM8K
Top 3 MSWs ZIP Top 3 MSWs ZIP Top 3 MSWs ZIP
0-CoT Let’s 50.13 Think 35.13 - -
Step-by-step 48.55 Let’s 30.33
Think 46.66 Step-by-step 22.05
0-CoTB Step-by-step 26.19 Take 81.61 Take 61.94
Breath 80.33 Deep 61.91
Step-by-step 80.19 Step-by-step 60.57
0-CoTR Work 58.60 Let’s 69.15 Let’s 51.11
Let’s 57.64 Work 69.00 Step-by-step 50.22
Step-by-step 65.72 Work 44.73
0-IRR Ignore 40.24 Feel 14.33 - -
Free 36.88 Ignore 13.03
Description 36.35 Free 12.88
0-PS Let’s 65.54 Solve 64.76 Problem 64.55
Problem 62.72 Problem 63.94 Let’s 61.50
First 62.30 Understand 59.57
WMT 19: German WMT 19: Chinese
Top 3 MSWs ZIP Top 3 MSWs ZIP
0-DSP Step-by-step 11.58 Please 11.15
Translation 9.94 Step-by-step 7.00
Complete 9.77 Translation 6.55
0-DTG Refine 13.32 Please 5.64
Firstly 12.90 Type 5.55
Translation 12.65 Firstly 5.53
Table 14: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on Llama 2, with the most significant word in bold. All reported words are confirmed as significantly important.
Prompt Type Correlation (r) Significance (p)
0-CoT -0.9948 0.064
0-CoTB -0.9954 0.060
0-CoTR -0.9982 0.037*
0-IRR -0.9999 0.005**
0-PS -0.9970 0.049*
Table 15: Pearson correlations between ZIP scores and GPT-4’s accuracy across classification tasks for different zero-shot prompts. All correlations show strong negative relationships (r < -0.99), with several reaching statistical significance (*p < 0.05, **p < 0.01). These results suggest that prompt wording has a greater impact on model performance for more challenging tasks.
A.4 Human Evaluation
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 11: Comparison of human judgments vs model-derived word importance for the CoTB prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 12: Comparison of human judgments vs model-derived word importance for the CoTR prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 13: Comparison of human judgments vs model-derived word importance for the IRR prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 14: Comparison of human judgments vs model-derived word importance for the PS prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 15: Comparison of human judgments vs model-derived word importance for the T1 prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 16: Comparison of human judgments vs model-derived word importance for the T2 prompt.
Figure 17: Google evaluation form for studying human intuition on word importance.
A.5 Qualitative Analysis
Task Information MSW LLM Answer with Original Prompt (Partial) Replaced Word LLM Answer with Perturbed Prompt (Partial)
Niles is collecting his book club’s annual fees. Each of the six members pays \$150/year towards snacks, plus \$30 each for six hardcover books and \$12 each for six paperback books. How much money does Niles collect in total?
True Answer: 2412
Prompt: 0-IRR
Dataset: GSM8k
Irrelevant 1. Each member pays an annual fee of \$150 for snacks.
2. Each member buys 6 hardcover books at \$30 each: 6×30 = 180
3. Each member buys 6 paperback books at \$12 each: 6×12 = 72
Now, we can calculate the total amount paid by each member: 150 + 180 + 72 = 402
Next, we find the total amount collected from all 6 members: 6×402 = 2412
The final answer is 2412.
Unrelated 1. Each member pays:
- Annual fee for snacks: \$150
- Cost for six hardcover books: \$30
- Cost for six paperback books: \$12
Total per member: 150 + 30 + 12 = 192
2. There are 6 members, so the total amount collected by Niles is: 192×6 = 1152
The final answer is 1152.
Alice, Bob, Claire, Dave, and Eve are playing a game. …. Finally, Dave and Alice swap balls. At the end of the game, Bob has the yellow ball. Options: (A) yellow ball. (B) white ball. (C) red ball. (D) pink ball. (E) blue ball.
True Answer: A
Prompt: 0-CoTR
Dataset: Big Bench
Right Bob has the yellow ball.
Correct answer is (A).
- Bob has the yellow ball.
Correct answer is (B).
A banana costs Rs.3 and an apple costs Rs.4. What will be the total cost of 4 dozen apples and bananas?
True Answer: 336
Prompt: 0-CoTB
Dataset: AQUA-RAT
step-by-step 1. Calculate the number of apples and bananas in 4 dozen:
- 1 dozen = 12
- Therefore, 4 dozen = 4 × 12 = 48 pieces of fruit.
2. Determine the cost of apples and bananas:
- The cost of one banana is Rs. 3.
- The cost of one apple is Rs. 4.
3. Calculate the total cost for 48 fruits:
- Let’s assume we buy 24 bananas and 24 apples (since we have no specific distribution given).
- Cost of 24 bananas = 24 × Rs. 3 = Rs. 72.
- Cost of 24 apples = 24 × Rs. 4 = Rs. 96.
4. Add the costs together:
- Total cost = Cost of bananas + Cost of apples
- Total cost = Rs. 72 + Rs. 96 = Rs. 168.
- 1. Cost of One Banana: Rs. 3
2. Cost of One Apple: Rs. 4
Next, we calculate the cost for 4 dozen (which is 48) of each fruit:
- Cost of 4 dozen Bananas:
Cost of 4 dozen bananas = 48 bananas × 3 Rs/banana = 144 Rs
- Cost of 4 dozen Apples:
Cost of 4 dozen apples = 48 apples × 4 Rs/apple = 192 Rs
Now, we can add the costs of the bananas and apples together to find the total cost:
Total cost = Cost of bananas + Cost of apples = 144 Rs + 192 Rs = 336 Rs
Table 16: Comparison of GPT-4o mini’s responses to original and perturbed prompts in classification tasks. Each row shows how changes to the most significant word (MSW) through synonym/co-hyponym replacement or removal affect model output.
Original Sentence True Translation Prompt Translation Language MSW LLM translation with original prompt Replaced word LLM translation with perturbed prompt
Der Vater unseres Landes wusste etwas von schlechter Presse. The father of our country knew something about bad press. 0-DTG German to English Refine The father of our nation knew something about bad press. Enhance The founding father of our nation was aware of the negative publicity.
 {CJK*}UTF8gbsn推动工作的新思路新举措新成效 To promote new ideas, new actions, and new effects of the work. 0-DPS Chinese to English Translation promote new ideas, new measures, and new results in work. Paraphrase advancing efforts with innovative concepts, new initiatives, and achieving new outcomes.
Table 17: Comparative analysis of original and perturbed prompt responses by GPT-4o mini on selected translation task instances where modifications to the most significant word (MSW) of the instructional prompt are made through synonym or co-homonym replacement.
Prompt: 0-IRR Dataset: GSM8k [Uncaptioned image]
Prompt: 0-CoTR Dataset: Big Bench [Uncaptioned image]
Prompt: 0-CoTB Dataset: AQUA-RAT [Uncaptioned image]
Table 18: The ZIP score heatmaps of the instructional prompts applied to the classification datasets featured in the qualitative analysis.
Prompt: 0-DTG Language: German to English [Uncaptioned image]
Prompt: 0-DPS Language: Chinese to English [Uncaptioned image]
Table 19: The ZIP score heatmaps of the instructional prompts applied to the translation datasets featured in the qualitative analysis.
A.6 Prompt Templates
Synonym Generation Prompts
[Ex: 1, Ex: 2, …, Ex: n] ← Few-shot examples applied for context.
Original Sentence: [Instructional Prompt]
Target word: [Target Word]
Task: Please provide 10 different meaningful alterations of the original sentence.
Each time replacing the word [Target Word] with a different synonym. Ensure the rest of the sentence remains unchanged.
Write down the altered sentence and the replaced word as the output.
Output:
Co-hyponym Generation Prompts
[Ex: 1, Ex: 2, …, Ex: n] ← Few-shot examples applied for context.
Original Sentence: [Instructional Prompt]
Target word: [Target Word]
Task: Please provide 10 different meaningful co-hyponyms of the original sentence.
Each time, replacing the word [Target Word] with a different co-hyponym. Ensure the rest of the sentence remains unchanged.
Write down the altered sentence and the replaced word as the output.
Output:
Meaningfulness and Correctness Prompts
Is this sentence meaningful and grammatically correct? “[Perturbed prompt]”
Answer only with Yes or No.
Table 20: Prompts used for creating perturbations. This task was preceded by a few-shot example set to guide the model in generating contextually relevant synonyms.
Classification-Based Tasks
Task Prompt template
GSM8k (0-CoT, 0-CoTB, 0-CoTR, 0-IRR, 0-PS) [Question]. [Instructional Prompt]
Write down your final answer to the question in this format: “The final answer is X.”
The type of X should be a number.
AQUA (0-CoT, 0-CoTB, 0-CoTR, 0-IRR, 0-PS) Multiple-Choice Question: [Question]
Answer Choices: [Options]
[Instructional Prompt]
Write down your final answer in the format: “The correct answer is [X].” where X is the letter of the correct answer choice (A, B, C, D, or E).
Big Bench (0-CoT, 0-CoTB, 0-CoTR, 0-IRR, 0-PS) Multiple-Choice Question: [Question]
Answer Choices: [Options]
[Instructional Prompt]
Write down your final answer in the format: “The correct answer is (X).” where X is the letter of the correct answer choice (A, B, C, D, or E).
Translation-Based Tasks
Task Prompt template
Translation (0-DSP) [Instructional Prompt]. [Original Sentence]
Before you write down the final English translation, please use these exact words: “####The final English translation of the complete sentence is:”
Translation (0-DTG) Step 1: Given the sentence: [Original Sentence], what is the English translation?
Before you write down the final English translation, please use these exact words: “####The final English translation of the complete sentence is:”
Step 2: Given the sentence: [Original Sentence], the English translation is [LLM Translation from Step 1].
[Instructional Prompt]. Before you write down the final English translation, please use these exact words: “####The final English translation of the complete sentence is:”
Table 21: Prompt templates used for both classification and translation tasks across various datasets.
A.7 Perturbations
Zero-shot
Instructional Prompt
Generated Candidates Semantic Similarity >30%
Meaningful and
grammatically correct
Final Perturbations
Let’s think step-by-step.
Let’s think slowly.
Let’s think. (removal)
Let’s think bit-by-bit.
Let’s think piecemeal.
Let’s think slowly.
Let’s think.
Let’s think bit-by-bit.
Let’s think piecemeal.
Let’s think.
Let’s think bit-by-bit.
Let’s think piecemeal.
Let’s think.
Let’s think bit-by-bit.
Take a deep breath and
work on this problem
step-by-step.
Take a deep breath and focus
on this problem step-by-step.
Take a deep breath and
on this problem step-by-step.
Take a deep breath and reflect
on this problem step-by-step.
Take a deep breath and study
on this problem step-by-step.
Take a deep breath and focus
on this problem step-by-step.
Take a deep breath and
on this problem step-by-step.
Take a deep breath and reflect
on this problem step-by-step.
Take a deep breath and study
on this problem step-by-step.
Take a deep breath and focus
on this problem step-by-step.
Take a deep breath and
on this problem step-by-step.
Take a deep breath and study
on this problem step-by-step.
Take a deep breath and focus
on this problem step-by-step.
Take a deep breath and study
on this problem step-by-step.
Let’s work this out in a
step-by-step way
to be sure we have the
right answer.
Let’s work this out in a
step-by-step way
to be secured we have the
right answer.
Let’s work this out in a
step-by-step way
to be we have the
right answer. (removal)
Let’s work this out in a
step-by-step way
to be concrete we have the
right answer.
Let’s work this out in a
step-by-step way
to be certain we have the
right answer.
Let’s work this out in a
step-by-step way
to be secured we have the
right answer.
Let’s work this out in a
step-by-step way
to be we have the
right answer.
Let’s work this out in a
step-by-step way
to be concrete we have the
right answer.
Let’s work this out in a
step-by-step way
to be certain we have the
right answer.
Let’s work this out in a
step-by-step way
to be we have the
right answer.
Let’s work this out in a
step-by-step way
to be certain we have the
right answer.
Let’s work this out in a
step-by-step way
to be certain we have the
right answer.
Feel free to ignore irrelevant
information in the problem
description.
Feel free to ignore insignificant
information in the problem
description.
Feel free to ignore
information in the problem
description. (removal)
Feel free to ignore irrelative
information in the problem
description.
Feel free to ignore unimportant
information in the problem
description.
Feel free to ignore insignificant
information in the problem
description.
Feel free to ignore
information in the problem
description.
Feel free to ignore irrelative
information in the problem
description.
Feel free to ignore unimportant
information in the problem
description.
Feel free to ignore insignificant
information in the problem
description.
Feel free to ignore information
in the problem description.
Feel free to ignore unimportant
information in the problem
description.
Feel free to ignore insignificant
information in the problem
description.
Feel free to ignore information
in the problem description.
Feel free to ignore unimportant
information in the problem
description.
Let’s first understand the
problem and devise a
plan to solve the problem.
Then let’s carry out the
plan and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to tackle the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to cope the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to crack the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to tackle the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to cope the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to crack the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to tackle the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to cope the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Let’s first understand the
problem and devise a plan
to tackle the problem.
Then let’s carry out the plan
and solve the problem
step-by-step.
Table 22: Example of perturbation filtering for classification prompts. Bold words perturbed via synonyms, co-hyponyms, removal. Filtered by semantic similarity (>30%) and grammar. Rejections in red.
Zero-shot
Instructional Prompt
Generated Candidates Semantic Similarity >30%
Meaningful and
grammatically correct
Final Perturbations
Please provide the translation
for the following sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the upcoming sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the after sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the next sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the consecutive sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the upcoming sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the after sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the next sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the consecutive sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the consecutive sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the consecutive sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the next sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the upcoming sentence
step-by-step and then provide
the complete sentence.
Please provide the translation
for the next sentence
step-by-step and then provide
the complete sentence.
Please detect the error type
firstly and refine the
translation then.
Please observe the error type
firstly and refine the
translation then.
Please the error type
firstly and refine the
translation then. (removal)
Please notice the error type
firstly and refine the
translation then.
Please discern the error type
firstly and refine
the translation then.
Please pick out the error type
firstly and refine
the translation then.
Please observe the error type
firstly and refine
the translation then.
Please the error type
firstly and refine the
translation then.
Please notice the error type
firstly and refine the
translation then.
Please discern the error type
firstly and refine the
translation then.
Please pick out the error type
firstly and refine the
translation then.
Please observe the error type
firstly and refine
the translation then.
Please the error type
firstly and refine the
translation then.
Please notice the error type
firstly and refine the
translation then.
Please discern the error type
firstly and refine the
translation then.
Please observe the error type
firstly and refine
the translation then.
Please notice the error type
firstly and refine the
translation then.
Please discern the error type
firstly and refine the
translation then.
Table 23: Example of perturbation filteringg process for translation task prompts. Bold words are perturbed via synonyms, co-hyponyms, and removal. Perturbations are filtered by semantic similarity (>30%) and grammatical correctness. Rejections in red.
Original Word Generated Candidates 20% 30% 40% 50%
Let’s We should (49.13%) Accept Accept Accept Reject
It is recommended that we (21.29%) Accept Reject Reject Reject
We can (51.57%) Accept Accept Accept Accept
First Before anything else (47.05%) Accept Accept Accept Reject
Right off the bat (25.66%) Accept Reject Reject Reject
Understand Perceive (58.85%) Accept Accept Accept Accept
Apprehend (42.68%) Accept Accept Accept Reject
Problem Dilemma (72.35%) Accept Accept Accept Accept
Hurdle (35.48%) Accept Accept Reject Reject
Difficulty (60.58%) Accept Accept Accept Accept
Devise Design (37.65%) Accept Accept Reject Reject
Draft (36.57%) Accept Accept Reject Reject
Set up (36.85%) Accept Accept Reject Reject
Plan Procedure (44.23%) Accept Accept Accept Reject
Strategy (57.06%) Accept Accept Accept Accept
Solve Crack (29.82%) Accept Reject Reject Reject
Tackle (40.75%) Accept Accept Accept Reject
Step-by-step Progressively (38.88%) Accept Accept Reject Reject
Phase by phase (30.38%) Accept Accept Reject Reject
Inch by inch (21.12%) Accept Reject Reject Reject
Accuracy 85% 95% 65% 40%
Table 24: Evaluation of semantic similarity thresholds (20%-50%) on the 0-PS prompt using 20 manually validated word variants. Each row shows an original word, its generated alternatives with similarity scores, and acceptance decisions across thresholds. Green indicates correct decisions (accept/reject), red indicates incorrect ones. The accuracy row demonstrates that 30% provides optimal performance (95%) in identifying meaningfully related perturbations.
A.8 Token Usage per Instructional Prompt
AQUA-RAT Big Bench GSM8K
Input tokens Output tokens Total tokens Input tokens Output tokens Total tokens Input tokens Output tokens Total tokens
0-CoT 600765 1825656 2426421 871710 1204340 2076050 488985 1098887 1587872
0-CoTB 1735827 3800599 5536426 2464368 2502721 4967089 1435263 3207869 4643132
0-CoTR 2472444 4488703 6961147 3447846 2974113 6421959 2070036 4367080 6437116
0-IRR 1390389 3014827 4405216 1974426 1831424 3805850 1149441 1895271 3044712
0-PS 4088076 9379794 13467870 5581284 6834640 12415924 3472044 7837170 11309214
Total 10287501 22509579 32797080 14339634 15347238 29686872 8615769 18406277 27022046
Table 25: Estimation of token usage per instructional prompt across different datasets (AQUA-RAT, Big Bench, and GSM8K) using GPT-4o mini. This table presents the input tokens, output tokens, and total tokens calculated with space-based tokenization. Actual tokenization methods employed by GPT-4o mini may vary, and thus these figures serve as approximations of token usage.