### v3

# Which Words Matter Most in Zero-Shot Prompts?

Abstract
While zero-shot instructional prompts like “Let’s think step-by-step” have revolutionized Large Language Model performance, a fundamental question remains unanswered: which specific words drive their remarkable effectiveness? We introduce the ZIP score (Zero-shot Importance of Perturbation), the first systematic method to quantify individual word importance in instructional prompts through controlled perturbations including synonym replacement, co-hyponym substitution, and strategic removal. Our analysis across four flagship models, seven widely-adopted prompts, and multiple task domains reveals four key findings: (1) Task-specific word hierarchies exist where mathematical problems prioritize “step-by-step” while reasoning tasks favor “think”; (2) Proprietary models show superior alignment with human intuitions compared to open-source alternatives; (3) Nouns dominate importance rankings, consistently representing the majority of significant words; and (4) Word importance inversely correlates with model performance, indicating prompts have greatest impact where models struggle most. Beyond revealing these patterns, we establish the first ground-truth benchmark for prompt interpretability through 20 validation prompts with predetermined key words, where ZIP achieves 90% accuracy versus LIME’s 60%. Our findings advance prompt science, the study of how language shapes model behavior, providing both practical insights for prompt engineering and theoretical understanding of word-level effects in LLMs. [^1] The complete codebase and documentation of language model interactions will be publicly available upon publication.
Which Words Matter Most in Zero-Shot Prompts?
Nikta Gohari Sadr1, Sangmitra Madhusudan1, Hassan Sajjad2 and Ali Emami3 1Brock University, St. Catharines, Canada 2Dalhousie University, Halifax, Canada 3Emory University, Atlanta, USA
1 Introduction
Figure 1: ZIP score generation process for the Chain-of-Thought prompt using GPT-4o mini on a AQUA-RAT dataset instance. Perturbed prompts are compared to the original, generating word-level ZIP scores. The red box highlights “step-by-step” as significantly important based on statistical analysis.
Chain-of-Thought prompting transformed how we use Large Language Models. Adding “Let’s think step-by-step” to a prompt, GPT-3 suddenly solved complex math problems it previously failed on Wei et al. (2022). This discovery sparked a wave of instruction prompts: Self-Consistency Wang et al. (2023b), Plan-and-Solve Wang et al. (2023a), and others, each improving performance through carefully crafted phrases. These prompts didn’t just boost task performance; they became fundamental to training Large Reasoning Models like OpenAI’s o3 OpenAI (2025).
Consider the Plan-and-Solve prompt: “Let’s first understand the problem and devise a plan to solve it. Then, solve it step-by-step” Wang et al. (2023a). This manually crafted extension of Chain-of-Thought consistently outperforms the original prompt across benchmarks. The intuition seems clear: explicit planning should reduce errors. But which words actually drive this improvement? Is it the instruction to “plan”? The directive to “understand”? Or does the familiar “step-by-step” do most of the work? Despite the empirical success of these prompts, the selection of which words to include remains largely ad hoc, we lack systematic understanding of how individual word choices affect model performance.
While it is well-known that models are sensitive to prompt structure Sclar et al. (2024), the importance of individual word choices in a prompt remains poorly understood. Individual word choices can have unexpected and counterintuitive effects on model performance. For instance, replacing “solve” with its synonym “work out” seems harmless, but can produce vastly different outputs.
Following established work on perturbation-based interpretability Choudhary et al. (2022); Saleem et al. (2022); Feng et al. (2018), we define word importance as the degree to which a word’s presence, absence, or modification impacts model performance on a given task. While gradient-based Wallace et al. (2019); Yin and Neubig (2022); Ferrando et al. (2023); Wu et al. (2023) and attention-based methods Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022) also aim to identify important words, they infer importance from model internals rather than directly measuring performance impact.
Beyond this, methods that analyze internal representations face practical limitations. They require access to model internals, making them inapplicable to closed-source models like GPT-4. They also face reliability issues: gradients can be manipulated Wang et al. (2020) and attention weights often fail to correlate with actual performance impact Jain and Wallace (2019); Ethayarajh and Jurafsky (2021). Recent work on mechanistic interpretability (i.e., circuit analysis, sparse autoencoders, and activation patching) Lindsey et al. (2025) provides detailed internal analysis but requires substantial computational resources and whitebox access. For understanding how word choices actually affect model outputs, we need lightweight, model-agnostic methods that directly measure performance changes.
We introduce the ZIP score (Zero-shot Importance of Perturbation score), a metric that quantifies word importance through systematic perturbations including synonym replacement, co-hyponym substitution, and word removal. Our approach works with any model, requiring only the ability to query the model. Additionally, we develop a set of validation prompts with known ground-truth word importance, providing a benchmark for evaluating interpretability methods. Figure 1 shows our method identifying “step-by-step” as the critical component in Chain-of-Thought prompting for a mathematical reasoning task.
Testing across four flagship models, seven prompts, and multiple tasks reveals surprising patterns. Word importance varies dramatically by task type: “step-by-step” dominates in mathematical problems while “think” matters more for common-sense reasoning. Proprietary models like GPT-4o mini show closer alignment with human intuitions than open-source alternatives. Across all models, nouns consistently emerge as most important (47.4%–65.9% of significant words). Most strikingly, ZIP scores inversely correlate with model performance $(|\text{r}|>0.9)$: prompts have their greatest impact on tasks where models struggle.
These findings advance prompt science Shah (2025), the systematic study of how prompts elicit model behavior, by quantifying word level importance in zero-shot instructional prompts. By revealing which words actually matter and how word choices affect model outputs, we provide both practical guidance for prompt design and deeper insights into prompt-model interactions.
2 Related Work
Feature importance interpretation in language models is typically approached through three main methods: gradient-based, attention-based, and perturbation-based Choudhary et al. (2022).
Gradient-Based Methods identify influential features by calculating the gradient of the output logits with respect to input elements Wallace et al. (2019). Recent advances include gradient-based post hoc contrastive explanations for model predictions Yin and Neubig (2022); Ferrando et al. (2023) and analyses of how CoT prompting affects saliency scores Wu et al. (2023). However, these methods face significant limitations: they require access to model internals, making them inapplicable to closed-source models. Furthermore, research has shown that model gradients can be easily manipulated, raising concerns about the reliability of gradient-based analyses Wang et al. (2020).
Attention-Based Methods interpret feature importance by analyzing the weighted sum of intermediate representations in neural networks Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022). While intuitive, these approaches have several drawbacks. For one, the attention weights do not always correspond to the most important features for model predictions Jain and Wallace (2019) and may not correlate well with other feature importance measures Ethayarajh and Jurafsky (2021). Moreover, attention weights often contain redundant and mixed information, leading to unreliable explanations Bastings and Filippova (2020). Like gradient-based methods, these approaches also require access to model internals, limiting their applicability to open-source models.
Perturbation-Based Methods can analyze any LLM regardless of architecture accessibility by measuring how outputs change when inputs are modified. As Choudhary et al. (2022) describes, “a word (token) or a collection of words (tokens) are modified or removed from the input samples, and a resulting change is measured.” Notable approaches include LIME Ribeiro et al. (2016), which creates local linear approximations of model behavior, and other dataset instance perturbations Zafar and Khan (2019). Recent research has extended these methods to few-shot demonstrations Liu et al. (2023) and system prompts Hackmann et al. (2024); Yin et al. (2023).
Despite these advancements, there remains a significant gap in the interpretation of zero-shot instructional prompts. Current research often uses basic token masking, which may oversimplify prompt semantics. It can also result in incoherent perturbations, potentially failing to interpret the true impact of the word on model predictions Yin et al. (2023); Feng et al. (2018). Our research addresses this by introducing a new metric that employs a variety of meaningful perturbations to reveal the importance of each word in zero-shot instructional prompts.
Prompt Science: Recent work distinguishes between prompt engineering (optimizing prompts for specific tasks) and prompt science (using prompts to discover regularities in model behavior) Holtzman and Tan (2025); Shah (2025). Studies have documented prompt brittleness, the sensitivity of model outputs to minor prompt variations Arora et al. (2022); Min et al. (2022); Sclar et al. (2023). For instance, Sclar et al. (2023) found that formatting changes alone can swing accuracy by over 70 percentage points. Rather than viewing this sensitivity as a flaw, we leverage it as a tool for understanding word importance. As Holtzman and Tan (2025) argued, prompt sensitivity reflects models’ attempts to “infer substantial information from limited context,” making it a feature for scientific investigation rather than a bug to be fixed. Our ZIP score provides a systematic framework for what they call “behavioral studies that use varied prompts in structured ways to confirm hypotheses.”
3 The ZIP score
The ZIP score quantifies the importance of individual words in a prompt by measuring how much the model’s performance changes when that word is perturbed or removed.
3.1 Formalization of the ZIP score
We formalize the ZIP score through the following steps, using a running example:
Example Task ($T$): Determine whether a number is prime or composite.
Original Prompt ($P$): “Take a deep breath and work on this problem step-by-step” Luo et al. (2024).
Step 1: Prompt Representation
Let $P$ represent the original prompt, where $P=w_{1},w_{2},\dots,w_{I}$ and $w_{i}$ denotes the $i$-th word. Throughout this paper, we use “word” to refer to any token resulting from our space-based tokenization method. This may include contractions, compound words, or multi-word expressions. Example:
$P=[\text{``Take''},\text{``a''},\text{``deep''},\ldots,\text{``step-by-step''}]$
Figure 2: Perturbation generation and filtering pipeline
Step 2: Perturbation Generation
For each word $w_{i}$ in $P$, we generate a set of perturbations $\{P_{i1},P_{i2},\dots,P_{iJ}\}$, through context-aware synonym replacement and co-hyponym replacement using GPT-4. Synonyms provide similar meanings, co-hyponyms introduce related concepts, and removal tests word necessity. Each $P_{ij}$ represents a modified version of $P$ where $w_{i}$ has been replaced or omitted.
Example: For $w_{4}=\text{``breath''}$:
  • $P_{41}=\text{``Take a deep pause and work on ...''}$
  • $P_{42}=\text{``Take a deep glimpse and work on ...''}$
  • $P_{43}=\text{``Take a deep look and work on ...''}$
Aligned with existing literature on perturbation-based methods Choudhary et al. (2022); Ivanovs et al. (2021); Saleem et al. (2022), we posit that a word’s importance is indicated by the model’s performance shifts in response to any meaningful alteration of that word, whether through subtle meaningful shifts or its complete removal (see §6 for analyses and demonstrations).
To create a final set of perturbations that provides insight into the impact of each word on the model output, as depicted in Figure 2, we filter the perturbed sentences as follows:
  1. 1.
    The Universal Sentence Encoder Cer et al. (2018) to measure semantic similarity between $P$ and $P_{ij}$, keeping those with >>30% similarity. This threshold was chosen to balance maintaining context while allowing meaningful alterations (see Table 21 in A.7).
  2. 2.
    GPT-4 evaluation to assess grammatical correctness and semantic viability of each $P_{ij}$.
Our filtering process excludes $P_{42}$ for insufficient semantic similarity and grammatical issues, leaving $P_{41}$ and $P_{43}$ as valid perturbations.
We maintained cost-effectiveness by designing concise prompts and compact output formats to minimize token usage, and by using the Universal Sentence Encoder to filter out unrelated perturbations before making API calls. Across the seven instructional prompts, this filtering process resulted in an average of 9.04 perturbed sentences per word. Detailed prompts and examples for perturbation creation and evaluation are provided in Appendix Sections A.6 and A.7.
Step 3: Model Prediction
We input the original prompt $P$ and each of its validated perturbed variants $P_{ij}$ into the language model $M$ to address an instance $(t)$ of task $T$.
Example ($t$): Is 29 prime or composite?
$\text{pred}_{t}^{M}(P)$ $=\text{``Prime''}\quad\text{(Correct)}$
$\text{pred}_{t}^{M}(P_{41})$ $=\text{``Prime''}\quad\text{(Correct)}$
$\text{pred}_{t}^{M}(P_{43})$ $=\text{``Composite''}\quad\text{(Incorrect)}$
Step 4: Disagreement Calculation
We define a disagreement function $d$ that measures the difference between model predictions for each perturbed prompt ($P_{ij}$) and the original prompt ($P$). This function varies by task type $T$. In our study, for classification tasks:
$d_{t}^{M}(P,P_{ij})=\begin{cases}1&\text{if }\text{pred}_{t}^{M}(P)\neq\text{pred}_{t}^{M}(P_{ij})\\ 0&\text{otherwise}\end{cases}$
For translation tasks:
$d_{t}^{M}(P,P_{ij})=|\text{BLEU}_{t}^{M}(P)-\text{BLEU}_{t}^{M}(P_{ij})|$
Here $\text{BLEU}_{t}^{M}(P)$ and $\text{BLEU}_{t}^{M}(P_{ij})$ are the BLEU scores Papineni et al. (2002) for translations generated from the original and perturbed prompts, respectively.
Since the above example $T$ is a classification task, we calculate disagreement scores as follows:
$d_{t}^{M}(P,P_{41})$ $=0\quad\text{(No disagreement)}$
$d_{t}^{M}(P,P_{43})$ $=1\quad\text{(Disagreement)}$
Step 5: ZIP score Calculation
Finally, for each word $w_{i}$, we compute $\text{ZIP}_{T}^{M}(w_{i})$ by averaging disagreement scores across all $J$ perturbations of $w_{i}$ and all dataset instances:
$\text{ZIP}_{T}^{M}(w_{i})=100\cdot\frac{1}{N}\sum_{t=1}^{N}\left(\frac{1}{J}\sum_{j=1}^{J}d_{t}^{M}(P,P_{ij})\right),$
where $N$ is the number of dataset instances, $P$ is the original prompt, and $P_{ij}$ is the $j$-th perturbation of word $w_{i}$ in the $t$-th dataset instance.
Example: For $w_{4}=\text{``breath''}$ in a single instance ($N=1$):
$\text{ZIP}_{T}^{M}(\text{``breath''})=100\cdot\frac{1}{2}(0+1)=50$
The ZIP score ($\in[0,100]$) represents the average percentage change in model performance observed across all tested dataset instances when perturbing a given word. This metric is adaptable to various performance measures, making it versatile for different types of tasks. Interpreting the ZIP score is straightforward: a higher score indicates greater word importance, as perturbing it leads to larger changes in model output. Conversely, a lower score suggests the word has less impact on performance for the given task.
Code Prompt Mutual Datasets
0-CoT Let’s think step-by-step. Kojima et al. (2024) BIG-bench, GSM8K, AQUA-RAT
0-CoTB Take a deep breath and work on this problem step-by-step. Yang et al. (2024) AQUA-RAT, GSM8K
0-CoTR Let’s work this out in a step-by-step way to be sure we have the right answer. Zhou et al. (2023) GSM8K, BIG-bench
0-IRR Feel free to ignore irrelevant information in the problem description. Shi et al. (2023) GSM8k
0-PS Let’s first understand the problem and devise a plan to solve it. Then, solve it step-by-step. (Plan & Solve Wang et al. (2023a)) GSM8K, AQUA-RAT
0-DSP Provide the translation step-by-step, then complete the sentence. Peng et al. (2023) WMT19 (German, Chinese)
0-DTG Detect the error type first, then refine the translation. (Deliberate then Generate Li et al. (2023)) WMT (German, Chinese)
Table 1: Zero-shot instructional prompts used in our experiments, and mutual datasets they were tested on.
3.2 Identifying Significantly Important Words
To determine which words are significantly important in prompts, we use a statistical approach that distinguishes between the inherent variability in model outputs and the effects of our perturbations:
  1. 1.
    For prompt word, we generate a set of perturbed prompts (the same set used in the ZIP score calculation).
  2. 2.
    We repeatedly input both the original prompt and each of its perturbed versions to the model, generating multiple outputs for each, up to the maximum number of word perturbations in the prompt.
  3. 3.
    We compare the ZIP scores from the original re-prompts with those from each perturbed prompt using the Wilcoxon rank-sum test.
The Wilcoxon rank-sum test tests whether two independent samples (i.e., the outputs from the original re-prompts and a perturbed prompt) come from the same distribution. We deem words with a p-value << 0.05 as ‘significantly important’.
4 Experimental Setup
4.1 Validation Prompts
Despite the success of zero-shot prompts, establishing a ground truth for word-level importance within them is challenging because we may not know which words should matter most. To address this, we created a set of 20 simple zero-shot instructional validation prompts, each designed so that a single predetermined ‘key word’ determines the correct output. For example, consider the prompt “Say the word green”. Here, “green” is the key word: if the model does not produce “green” exactly, it fails. By knowing in advance which word is crucial for a correct output, we can objectively assess the effectiveness of interpretability methods.
We constructed 20 validation prompts (Table 3) with varied syntax and structure, such as “Print the digits 123, “Type the letter X, and “Print carrot with no additional text”. Each prompt contains a key word that directly determines the correct output, any alteration to this word should change the model’s response. For example, in “Print the digits 123”, only perturbations to “123” should affect whether the model outputs “123” correctly.
4.2 Prompts
We evaluated seven instructional prompts across various datasets, as detailed in Table 1. These prompts were selected based on their proven effectiveness and generalization capabilities in recent literature. We used a simple space-based tokenization method, where tokens are defined as sequences of characters separated by whitespace. This approach preserves contractions and compound words (e.g., “Let’s” as one token), which are often semantically important in instructional prompts. Dataset-specific prompts are included in the Appendix A.6 (Table 19).
4.3 Datasets
We selected a diverse range of datasets known for effectively testing zero-shot prompts, as highlighted in bold in Table 1. For classification tasks, datasets included GSM8K Cobbe et al. (2021), AQUA-RAT Ling et al. (2017), and the Big Bench dataset Srivastava et al. (2023). Translation tasks used the WMT19 dataset, concentrating on Chinese and German languages. We evaluated on 150 instances per dataset, which we determined was sufficient to provide statistically significant results while managing computational costs.
4.4 Models
Four models were used in our experiments: GPT-4o mini, GPT-3.5-turbo, Llama-2-70B-chat, and Mixtral-8x7B-Instruct-v0.1 OpenAI (2023); Touvron et al. (2023); Jiang et al. (2024). Each model was configured with a temperature of 0.5 to balance between deterministic outputs and allowing for some variability. For token usage estimates per prompt, see Appendix A.8.
4.5 Baselines
We compare our ZIP method with LIME Ribeiro et al. (2016), a widely-used perturbation-based explanation method. For validation prompts, we configured LIME to treat the task as binary classification: Label 1 represents the target word (e.g., “green” for “Say the word green”), and Label 2 represents any other output. LIME automatically creates perturbations through random word removal and fits a linear surrogate model to quantify word importance, using 50 perturbations per instance.
4.6 Human Evaluation
We conducted a study with 15 diverse participants proficient in English to compare human and model interpretations of the instructional language. For each prompt, participants selected up to three words that they deemed most important to solve a certain task. The evaluation form is shown in Appendix A.5 Fig. 17. We used the Jaccard Index to measure inter-annotator agreement across participants, yielding an average of 0.4714, indicating moderate agreement. Jaccard Index was chosen over Fleiss’ Kappa to accommodate multiple word selections per participant. Table 17 in the Appendix illustrates the selected words across participants and prompts. While there is moderate agreement overall, the table highlights clear variation in which words individuals judged as most important. This variation points at the subjectivity of human intuition and the need for model-based evaluations.
5 Results
5.1 Validation Results
Table 3 compares how ZIP and LIME performed in identifying key words on our 20 validation prompts on GPT-4o mini. ZIP correctly identified the predetermined key word in 90% of prompts, significantly outperforming LIME’s 60% accuracy. LIME’s errors often involved misidentifying contextual words (e.g., “Print,” “Output,” “Say") as most important, while ZIP maintained focus on the target words that directly determine the output.
ZIP’s strong performance was consistent across all four models (e.g., with 100% accuracy on GPT-3.5-turbo), validating ZIP’s effectiveness in identifying significant words in instructional prompts. Detailed results are provided in Appendix A.2.
5.2 Cross-Model Analysis
Table 2 presents ZIP score heatmaps for the CoT prompt “Let’s Think Step-by-Step” across models on Big Bench. Our analysis reveals:
  • “Step-by-Step” vs “Think”: While both words are important, step-by-step shows, on average, a higher ZIP score across models. This suggests that the explicit instruction for a structured approach (“step-by-step”) may be more universally interpreted and utilized by models than the more general cognitive instruction (“think”).
  • Proprietary vs Open-source models: Proprietary models (GPT-4o mini, GPT-3.5-turbo) identify fewer words as significantly important (1 word) compared to open-source models (Mixtral, Llama 2, with 2-3 words). This suggests that proprietary models are relatively robust, showing less dependence on single words in prompts, while open-source models exhibit broader sensitivity to individual word choices.
Results on all datasets are provided in Appendix A.1 Table 5.
GPT-4o mini [Uncaptioned image]
GPT 3.5 Turbo [Uncaptioned image]
Mixtral [Uncaptioned image]
Llama 2 [Uncaptioned image]
Table 2: ZIP score heatmaps for the CoT prompt across four models on the Big Bench dataset. Red boxes indicate significantly important words.
Validation Prompts ZIP LIME
Say the word green Green Green
Print the digits 123 123 Print
Output the color blue Blue Output
Print carrot with no additional text Print Carrot
Display the word circle Circle Circle
When you’re ready just say coffee Coffee Coffee
Return the value five Five Five
Respond with only the word hello Hello Hello
Repeat the term mirror Mirror Repeat
Nine is the number you should write Nine Nine
Answer with the word pizza Pizza With
Begin by writing hello and then finish Hello Hello
Type purple in your response now Purple Purple
Say red and then stop Red Say
When you respond lead with river River River
Write the number seven Seven Number
Carefully type silver when responding here Silver Silver
Enter the word tomato Tomato The
Can you mention the direction up in our chat? Direction Chat
Type the letter X X X
Accuracy 90% 60%
Table 3: Most important words as identified by ZIP and LIME for GPT-4o mini on validation prompts. Underlined words are the target keywords in the prompt. Bold indicates failure to identify the correct keyword.
5.3 Task-Specific Patterns
AQUA-RAT Big Bench GSM8K
Top 3 MSWs ZIP Top 3 MSWs ZIP Top 3 MSWs ZIP
0-CoT Step-by-step 28.44 Think 3.80 Step-by-step 5.83
0-CoTB Step-by-step 34.49 Problem 1.95 Step-by-step 6.57
Work 1.33
0-CoTR Answer 27.66 Right 4.66 Answer 5.33
Sure 27.61 Work 4.40 Step-by-step 5.22
Right 26.44 Sure 4.09 Way 5.21
0-IRR Description 30.71 Ignore 3.87 Irrelevant 6.51
Ignore 30.42 Description 3.74 Description 6.41
Irrelevant 30.30
0-PS Plan 28.50 Plan 4.05 Step-by-step 7.28
Step-by-step 28.40 First 3.94 Problem 6.83
Solve 27.72 Solve 3.72 Solve 6.54
WMT 19: German WMT 19: Chinese
Top 3 MSWs ZIP Top 3 MSWs ZIP
0-DSP Translation 10.92 Translation 6.57
Step-by-step 7.68 Step-by-step 5.50
Sentence 6.50 Sentence 5.35
0-DTG Refine 8.43 Refine 5.31
Error 8.33 Type 4.92
Detect 8.06 Error 4.92
Table 4: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on GPT-4o Mini (most significant in bold). All reported words are confirmed as significantly important.
Table 4 presents the ZIP score results for various zero-shot instructional prompts (detailed in Table 1) using the GPT-4o mini model. Our analysis demonstrates distinct patterns in the importance of words across various datasets and task types:
  • For mathematical and algebraic tasks (AQUA-RAT and GSM8K), the phrase step-by-step consistently emerges as highly important.
  • In contrast, for common sense tasks (Big Bench), words like think and problem show higher importance.
  • For translation tasks, translation and refine are identified as the most important words, which demonstrates the significance of task-specific instructions in multilingual contexts.
  • ZIP scores strongly correlate inversely with model performance $(\lvert r\rvert>0.9)$ across all prompts. For example, using CoT, GPT-4o mini shows higher ZIP scores on AQUA-RAT (accuracy: 68.26%, ZIP: 28.44) compared to GSM8K (91.59%, 5.83) and Big Bench (96.93%, 3.80) indicating that instructional prompts have greater impact on challenging tasks (Table 15).
Results for other models (similar trends) are provided in Appendix A.3 Tables 12 - 14.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 4: Part-of-Speech (POS) distribution of significantly important words across four LLMs.
5.4 Part-of-Speech Analysis
Figure 4 shows the part-of-speech distribution of the most important words in the seven prompts. Key observations are:
  • Noun dominance: Nouns (NN) consistently rank as most important across all models (47.4%–65.9% of significant words).
  • Model patterns: GPT-4o mini and GPT-3.5-turbo, sharing similar foundational architectures despite vastly different parameter sizes, show similar POS distributions; Mixtral and Llama 2 show more balanced POS importance.
  • Verb importance: Base form verbs (VB) consistently rank 2nd in importance across all models.
  • Inter-model variation: Adverbs (RB) and non-3rd person singular present verbs (VBP) show the highest variation in importance across models.
5.5 Human Alignment
Figure 5 presents radar plots comparing human and model-derived word importance for the CoT prompt. Human values represent the proportion of evaluators selecting each word as important, while model values show the proportion of tasks where each word was significantly important.
Proprietary models (GPT-4o mini and GPT-3.5-turbo) align strongly with human judgments, while open-source models show less alignment, particularly Llama 2. “Step-by-step” emerges as highly important across all evaluations, while “Let’s” is consistently less emphasized. Similar analyses for remaining prompts are provided in Appendix A.5 Figures 11 - 16.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 5: Comparison of human judgments vs model-derived word importance for the CoT prompt.
6 Qualitative Analysis
Table 16 shows that small changes to prompt wording can lead to substantial differences in model output, highlighting the significance of individual words in shaping model reasoning. For example, the model correctly solved the museum ticket problem under the original 0-CoT prompt. However, removing “step-by-step” reduced its reasoning from five steps to two and led it to incorrectly apply the discount to the grandfather, an assumption that may seem reasonable but is not stated in the question. In a more counterintuitive case, removing “step-by-step” actually improved performance by preventing the model from introducing a faulty assumption. Other substitutions also showed significant effects. Replacing “solve” with “work out” in the 0-PS prompt disrupted arithmetic reasoning, and swapping “irrelevant” with “unrelated” in the IRR prompt caused the model to overlook numbers expressed in words, halving the final answer.
These examples reveal that seemingly equivalent words can produce significantly different model outputs. This observation supports our dual approach of using both synonym/co-hyponym substitution as well as word removal to reveal different aspects of how models interpret prompt wording.
7 Conclusion
This study introduces the ZIP score, a novel metric for quantifying word importance in zero-shot instructional prompts. Our findings reveal interesting patterns in word importance across models and tasks. For instance, while both “step-by-step” and “think” show high ZIP scores, which one is more influential depends on the model and task. We observed significant differences in word importance and alignment with human intuition between proprietary and open-source models. These insights aid our understanding of LLM behavior and open new avenues for prompt engineering and model analysis. Future work could explore the use of these findings to design more effective prompts tailored to specific model architectures.
Limitations
Perturbation Design Choices: Our study used three well-established perturbation types (synonym replacement, co-hyponym substitution, and removal) that align with established interpretability literature Choudhary et al. (2022). This focused approach allowed systematic comparison and validation against human judgments. While additional perturbation types do exist, our chosen methods successfully captured the key patterns of word importance, as demonstrated by our superior performance compared to existing baselines.
Individual Word Analysis Framework: Following standard practice in interpretability research, we analyze words independently to allow for clear baselines for word-level importance. This approach, while necessarily simplified, provides the foundational understanding required before investigating more complex word interactions, which we hope will be the basis behind a vast new line of works. Our strong alignment with human judgments, particularly in proprietary models, validates this methodology as an effective starting point for prompt analysis.
Model and Task Coverage: We selected four architecturally diverse models representing both proprietary (GPT-4o mini, GPT-3.5-turbo) and open-source (Mixtral, Llama-2) approaches, covering the major LLM families in current use. Our goal is not to demonstrate universality across every possible model (a nice-to-have, nonetheless!), but rather to establish that consistent word-level importance patterns exist within the models we study. Our focus was on spanning several classification and translation domains, and showing the method’s applicability across different output types and evaluation metrics.
Computational Efficiency Considerations: Our sample size of 150 instances per dataset was chosen to ensure statistical significance while maintaining computational tractability for systematic perturbation analysis. This approach successfully identified clear and consistent patterns of word importance across models and tasks with appropriate statistical rigor. Of course, larger sample sizes could provide additional statistical power, though our current approach demonstrates sufficient precision and statistical significance for the specific research questions addressed.
Research Scope and Applications: This work establishes ZIP as a foundational tool for understanding prompt interpretability rather than prompt optimization. Our focus on analysis rather than engineering represents a deliberate choice to build scientific understanding first. The insights generated provide a principled foundation for future prompt engineering efforts. Please refer to (Holtzman and Tan, 2025) for a fascinating discussion on the distinction between prompt science and prompt engineering.
Ethical Considerations in Application of Interpretability Insights: Although our method offers a deeper understanding of LLM operations, it is important to consider the potential for misuse. Detailed knowledge of how LLMs weigh different prompt elements could potentially be used to tailor content in ways that could manipulate or bias decision-making processes. It is crucial to apply these insights ethically with a commitment to fairness.
Appendix A Appendix
A.1 Chain-of-Thought ZIP score Heatmaps
GPT-4o mini Refer to caption
GPT 3.5 Turbo Refer to caption
Mixtral Refer to caption
Llama 2 Refer to caption
GPT-4o mini Refer to caption
GPT 3.5 Turbo Refer to caption
Mixtral Refer to caption
Llama 2 Refer to caption
Table 5: ZIP score heatmaps of the Chain-of-Thought (CoT) prompt across four models using the AQUA-RAT and GSM8K datasets. The tables display the significant word identification performance for each model with different datasets. Red boxes highlight words identified as significantly important, confirmed through statistical analysis.
A.2 Controlled Pre-Experiment
Validation Prompts ZIP LIME
Say the word green. Green Green
Print the digits 123. 123 123
Output the color blue. Blue Color
Print carrot with no additional text. Print/Carrot Carrot
Display the word circle. Cricle Word
When you’re ready just say coffee. Coffee Say
Return the value five. Five Value
Respond with only the word hello. Hello The
Repeat the term mirror. Mirror Mirror/Repeat
Nine is the number you should write. Nine Nine
Answer with the word pizza. Word/Pizza With
Begin by writing hello and then finish. Hello Hello
Type purple in your response now. Purple Purple
Say red and then stop. Red Red
When you respond lead with river. River Lead
Write the number seven. Seven Number
Carefully type silver when responding here. Silver Silver
Enter the word tomato. Word/Tomato The
Can you mention the direction up in our chat? Up Up
Type the letter X. X X
Accuracy 100% 55%
Table 6: Most important words as identified by ZIP and LIME for GPT 3.5 Turbo on validation prompts. Underlined words are the target keywords in the prompt. Bold indicates failure to identify the correct key word.
Validation Prompts ZIP LIME
Say the word Green. Green Green
Print the digits 123. 123 Digits
Output the color blue. Color/Blue Blue
Print carrot with no additional text. Carrot Carrot
Display the word circle. Circle Circle
When you’re ready just say coffee. Coffee Coffee
Return the value five. Five Five
Respond with only the word Hello. Hello Hello
Repeat the term mirror. Mirror Term
Nine is the number you should write. Nine Nine
Answer with the word pizza. Pizza Pizza
Begin by writing hello and then finish. Hello Hello
Type purple in your response now. Purple Purple
Say red and then stop. Red Red
When you respond lead with river. River Respond
Write the number seven. Seven Seven
Carefully type silver when responding here. Silver Silver
Enter the word tomato. Tomato Tomato
Can you mention the direction up in our chat? Direction Chat
Type the letter X. X X
Accuracy 95% 80%
Table 7: Most important words as identified by ZIP and LIME for Mixtral on validation prompts. Underlined words are the target keywords in the prompt. Bold indicates failure to identify the correct key word.
Validation Prompts ZIP LIME
Say the word Green. Green Green
Print the digits 123. Print/Digits/123 Digits
Output the color blue. Blue Blue
Print carrot with no additional text. Carrot No
Display the word circle. Circle Display
When you’re ready just say coffee. Coffee Say
Return the value five. Five Value
Respond with only the word Hello. Hello Hello
Repeat the term mirror. Mirror Mirror
Nine is the number you should write. Nine Should
Answer with the word pizza. Pizza Pizza
Begin by writing hello and then finish. Writing Hello
Type purple in your response now. Purple Purple
Say red and then stop. Red Red
When you respond lead with river. River River
Write the number seven. Seven Number
Carefully type silver when responding here. Silver Here
Enter the word tomato. Tomato Tomato
Can you mention the direction up in our chat? Chat Direction
Type the letter X. X X
Accuracy 90% 55%
Table 8: Most important words as identified by ZIP and LIME for Llama 2 on validation prompts. Underlined words are the target keywords in the prompt. Bold indicates failure to identify the correct key word.
GPT-4o mini [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GPT 3.5 Turbo [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Llama 2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Mixtral [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 9: Heatmap visualization of LIME for five validation prompts on GPT-4o mini, GPT 3.5 Turbo, Llama 2, and Mixtral models. The color blue represents alignment with Label 1 (the desired target word), and orange indicates Label 2 (the second most probable alternative output), with the intensity of each color indicating the level of importance attributed by LIME. Check (✓) and cross (✗) marks indicate whether the model correctly identified the predefined most important word.
GPT-4o mini [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GPT 3.5 Turbo [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Llama 2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Mixtral [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 10: Heatmap visualization of ZIP method results for five validation prompts, applied to GPT-4o mini, GPT 3.5 Turbo, Llama 2, and Mixtral models. These heatmaps illustrate the method’s ability to identify the most significant words across different prompts. Red boxes indicate significantly important words. Check (✓) and cross (✗) marks indicate whether the ZIP method correctly identified the predefined most important word.
Prompt Identified MSW Model Output
Print the digits 123. 123 ✓ 123
Display the digits 123. 123
Print the digits _. 1234567890
Print the numbers 123. 123
Say the word green. Green ✓ Green
Say the word blue. Blue
Say the term green. Green
Say the word _. Hello
Repeat the term mirror. Mirror ✓ Mirror
Recite the term mirror. Mirror
Repeat the term glass. Glass
Repeat the word mirror. Mirror
Display the word circle. Circle ✓ Circle
Display the letter circle. O
Display the word Square. square
Display the word Round. round
Prompt Identified MSW Model Output
Print the digits 123 Print ✗ 123
_ the digits 123. One
_ _ digits _. Numbers
Print _ _ 123 123
Say the word green. Green ✓ Green
_ _ _ green Echo
Say the _ green Green
Say _ _ green Green
Repeat the term mirror. Repeat ✗ Mirror
_ the term mirror. Reflection
_ _ _ mirror. Reflect
Repeat _ term mirror. Mirror
Display the word circle. Circle ✓ Circle
_ _ word circle. Round
Display the word _. Do
Display _ word circle. Circle
Table 11: Comparison of GPT-4o Mini’s responses to the original validation prompt (bolded) and perturbed examples (underlined alterations), using ZIP and LIME methods. Check (✓) and cross (✗) marks indicate if the Most Significant Word (MSW) was identified correctly by each method.
A.3 ZIP Scores
AQUA-RAT Big Bench GSM8K
Top 3 MSWs ZIP Top 3 MSWs ZIP Top 3 MSWs ZIP
0-CoT Step-by-step 43.94 Step-by-step 48.38 Step-by-step 18.50
0-CoTB Problem 38.09 Step-by-step 38.71 Step-by-step 16.23
Breath 37.52 Breath 38.57
Problem 38.33
0-CoTR Sure 41.04 Right 39.59 Step-by-step 16.16
Sure 39.52
Let’s 38.88
0-IRR Feel 56.00 Description 49.43 Irrelevant 39.74
Description 50.15 Information 48.51 Ignore 38.06
Free 50.00 Free 48.22 Problem 35.09
0-PS Let’s 38.72 Plan 45.17 Devise 19.73
Carry 38.33 Problem 45.16 Plan 19.38
Problem 38.33 Let’s 45.08
WMT 19: German WMT 19: Chinese
Top 3 MSWs ZIP Top 3 MSWs ZIP
0-DSP Step-by-step 11.24 Step-by-step 7.90
Translation 8.78 Translation 6.92
Provide 7.39 Sentence 6.48
0-DTG Firstly 10.78 Firstly 7.34
Detect 10.04 Error 6.46
Translation 9.59 Detect 6.10
Table 12: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on GPT 3.5 Turbo, with the most significant word in bold. All reported words are confirmed as significantly important.
AQUA-RAT Big Bench GSM8K
Top 3 MSWs ZIP Top 3 MSWs ZIP Top 3 MSWs ZIP
0-CoT Let’s 58.53 Let’s 58.60 - -
Step-by-step 54.94 Step-by-step 58.50
0-CoTB Breath 58.38 Step-by-step 85.09 Step-by-step 33.66
Problem 57.47 Work 57.11 Breath 33.23
Take 57.11 Breath 54.66 Work 32.66
0-CoTR Step-by-step 55.22 Let’s 57.24 Step-by-step 32.16
Let’s 53.73 Work 57.00 Let’s 31.95
Way 50.42 Step-by-step 55.38 Work 29.93
0-IRR Ignore 46.48 Ignore 24.66 Description 45.12
Free 45.55 Irrelevant 21.23 Ignore 43.93
Description 42.92 Information 20.76 Irrelevant 43.02
0-PS Let’s 58.63 Carry 60.66 Understand 40.23
Solve 56.87 Understand 60.38 First 39.83
First 56.61 First 60.27 Problem 39.22
WMT 19: German WMT 19: Chinese
Top 3 MSWs ZIP Top 3 MSWs ZIP
0-DSP Translation 13.39 Step-by-step 7.55
Step-by-step 12.25 Translation 2.73
Following 12.02 Provide 2.34
0-DTG Detect 13.59 Please 8.00
Please 13.28 Type 7.91
Firstly 13.01 Firstly 7.89
Table 13: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on Mixtral, with the most significant word in bold. All reported words are confirmed as significantly important.
AQUA-RAT Big Bench GSM8K
Top 3 MSWs ZIP Top 3 MSWs ZIP Top 3 MSWs ZIP
0-CoT Let’s 50.13 Think 35.13 - -
Step-by-step 48.55 Let’s 30.33
Think 46.66 Step-by-step 22.05
0-CoTB Step-by-step 26.19 Take 81.61 Take 61.94
Breath 80.33 Deep 61.91
Step-by-step 80.19 Step-by-step 60.57
0-CoTR Work 58.60 Let’s 69.15 Let’s 51.11
Let’s 57.64 Work 69.00 Step-by-step 50.22
Step-by-step 65.72 Work 44.73
0-IRR Ignore 40.24 Feel 14.33 - -
Free 36.88 Ignore 13.03
Description 36.35 Free 12.88
0-PS Let’s 65.54 Solve 64.76 Problem 64.55
Problem 62.72 Problem 63.94 Let’s 61.50
First 62.30 Understand 59.57
WMT 19: German WMT 19: Chinese
Top 3 MSWs ZIP Top 3 MSWs ZIP
0-DSP Step-by-step 11.58 Please 11.15
Translation 9.94 Step-by-step 7.00
Complete 9.77 Translation 6.55
0-DTG Refine 13.32 Please 5.64
Firstly 12.90 Type 5.55
Translation 12.65 Firstly 5.53
Table 14: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on Llama 2, with the most significant word in bold. All reported words are confirmed as significantly important.
Prompt Type Correlation (r) Significance (p)
0-CoT -0.9948 0.064
0-CoTB -0.9954 0.060
0-CoTR -0.9982 0.037*
0-IRR -0.9999 0.005**
0-PS -0.9970 0.049*
Table 15: Pearson correlations between ZIP scores and GPT-4’s accuracy across classification tasks for different zero-shot prompts. All correlations show strong negative relationships (r << -0.99), with several reaching statistical significance (*p << 0.05, **p << 0.01). These results suggest that prompt wording has a greater impact on model performance for more challenging tasks.
A.4 Qualitative Analysis
Task Information LLM Answer (Original Prompt) Perturbation LLM Answer (Perturbed Prompt)
Dorothy is 15 years old and wants to go to a museum with her family. Her family consists of her, her younger brother, her parents, and her grandfather. The regular ticket cost is \$10. People 18 years old or younger have a discount of 30%. How much money will Dorothy have after this trip, when she currently has \$70?
Correct Answer: 26
Prompt: 0-CoT
Dataset: GSM8K
1. Identify family members:
2. Determine ticket prices:
- Regular ticket cost: \$10
- Discounted ticket price: \$10 - \$3 = \$7
3. Calculate the number of tickets:
- Dorothy (15 years old): 1 ticket at \$7
- Younger brother (10 years old): 1 ticket at \$7
- Parents (2 adults): 2 tickets at \$10 (= \$20)
- Grandfather (1 adult): 1 ticket at \$10
4. Calculate total cost: … \$44
5. Calculate money left after the trip:
\$70 - \$44 = \$26
Step-by-step $\downarrow$ (Removed) There are 3 people eligible for the discounted ticket (Dorothy, her brother, and her grandfather) and 2 adults who will pay full price. 1. Discounted tickets for Dorothy, her brother, and her grandfather: 3 $\times$ 7 = 21 2. Regular tickets for her parents: 2 $\times$ 10 = 20 Now, we sum the total cost of all tickets: 21 + 20 = 41 Next, we subtract the total ticket cost from Dorothy’s current amount of money: 70 - 41 = 29
A banana costs Rs. 3 and an apple costs Rs. 4. What will be the total cost of 4 dozen apples and bananas?
Correct Answer: 336
Prompt: 0-CoTB
Dataset: AQUA-RAT
1. Calculate the number of apples and bananas in 4 dozen: …
- Therefore, 4 dozen: 4 $\times$ 12 = 48 pieces of fruit.
2. Determine the cost of apples and bananas: …
3. Calculate the total cost for 48 fruits:
Let’s assume we buy 24 bananas and 24 apples (since we have no specific distribution given).
Cost of 24 bananas: 24 $\times$ 3 = 72 Rs.
Cost of 24 apples: 24 $\times$ 4 = 96 Rs.
4. Add the costs together:
Total cost: 72 + 96 = 168 Rs.
Step-by-step $\downarrow$ (Removed) 1. Cost of One Banana: Rs. 3
2. Cost of One Apple: Rs. 4
Next, we calculate the cost for 4 dozen (which is 48) of each fruit:
- Cost of 4 dozen Bananas: 48 $\times$ 3 = 144 Rs.
- Cost of 4 dozen Apples: 48 $\times$ 4 = 192 Rs.
Now, we can add the costs of the bananas and apples together to find the total cost:
144 + 192 = 336 Rs.
Mr. X died when his age was 1/59th of year of birth. What was his date of birth?
[’A)1947’, ’B)1948’, ’C)1980’, ’D)2000’, ’E)1998’]
Correct Answer: 1947
Prompt: 0-PS
Dataset: AQUA-RAT
1. Understanding the relationship: … This means if he was born in year $Y$ and died in year $D$, then his age at death can also be expressed as DYD-Y.
2. Setting up the equation: $D-Y=\frac{1}{59}Y$
3. Calculating the year of death: Since $D$ must be a whole number, $Y$ must be a multiple of 59.
4. Finding candidates for year of birth: The answer choices are: 1947, 1948, 1980, 2000, 1998.
5. Calculating multiples of 59:
- The multiples of 59 around the years given are: $59\times 33=1947$
Thus, the only candidate from our choices that is a multiple of 59 is 1947.
Solve $\downarrow$ Work out 1. Understanding the relationship between age and year of birth: … If he was born in year $Y$ and died in year $D$, his age at death can be expressed as DYD-Y.
2. Setting up the equation: $D-Y=\frac{1}{59}Y$
3. Finding possible years of birth: Since $D$ must be a whole number, $Y$ must be a multiple of 59.
4. Checking the answer choices:
- A: 1947: 1947/59 = 33 (not a whole number)
5. Conclusion: None of the given years are multiples of 59.
Niles is collecting his book club’s annual fees. Each of the six members pays \$150/year towards snacks, plus \$30 each for six hardcover books and \$12 each for six paperback books. How much money does Niles collect in total?
True Answer: 2412
Prompt: 0-IRR
Dataset: GSM8k
1. Each member pays an annual fee of \$150 for snacks.
2. Each member buys 6 hardcover books at \$30 each: 6 $\times$ 30 = 180
3. Each member buys 6 paperback books at \$12 each: 6 $\times$ 12 = 72
Now, we can calculate the total amount paid by each member: 150 + 180 + 72 = 402
Next, we find the total amount collected from all 6 members: 6 $\times$ 402 = 2412
The final answer is 2412.
Irrelevant $\downarrow$ Unrelated 1. Each member pays:
- Annual fee for snacks: \$150
- Cost for six hardcover books: \$30
- Cost for six paperback books: \$12
Total per member: 150 + 30 + 12 = 192
2. There are 6 members, so the total amount collected from Niles is: 192 $\times$ 6 = 1152
The final answer is 1152.
Table 16: Comparison of GPT-4o mini’s partial responses to the original zero-shot prompt and a perturbed version where one of the top 3 most important words identified by ZIP was modified.
A.5 Human Evaluation
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 11: Comparison of human judgments vs model-derived word importance for the CoTB prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 12: Comparison of human judgments vs model-derived word importance for the CoTR prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 13: Comparison of human judgments vs model-derived word importance for the IRR prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 14: Comparison of human judgments vs model-derived word importance for the PS prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 15: Comparison of human judgments vs model-derived word importance for the T1 prompt.
(a) GPT-4o mini
(b) GPT-3.5-turbo
(c) Mixtral
(d) Llama 2
Figure 16: Comparison of human judgments vs model-derived word importance for the T2 prompt.
0-CoT 0-CoTB 0-CoTR 0-IRR 0-PS
Participant 1
Think
step-by-step
Work
Problem
step-by-step
Work
Ignore
Irrelevant
Information
Understand
Problem
Plan
Participant 2
step-by-step
step-by-step
step-by-step
Right
Answer
Ignore
Irrelevant
Devise
Plan
step-by-step
Participant 3
Think
Work
Right
Irrelevant
Solve
Participant 4
step-by-step
step-by-step
step-by-step
Irrelevant
step-by-step
Participant 5
step-by-step
Breath
step-by-step
step-by-step
Sure
Right
Ignore
Irrelevant
Description
Understand
Plan
Solve
Table 17: Important words identified by participants across five zero-shot prompts for classification tasks.
Figure 17: Google evaluation form for studying human intuition on word importance.
A.6 Prompt Templates
Synonym Generation Prompts
[Ex: 1, Ex: 2, …, Ex: n] ← Few-shot examples applied for context.
Original Sentence: [Instructional Prompt]
Target word: [Target Word]
Task: Please provide 10 different meaningful alterations of the original sentence.
Each time replacing the word [Target Word] with a different synonym. Ensure the rest of the sentence remains unchanged.
Write down the altered sentence and the replaced word as the output.
Output:
Co-hyponym Generation Prompts
[Ex: 1, Ex: 2, …, Ex: n] ← Few-shot examples applied for context.
Original Sentence: [Instructional Prompt]
Target word: [Target Word]
Task: Please provide 10 different meaningful co-hyponyms of the original sentence.
Each time, replacing the word [Target Word] with a different co-hyponym. Ensure the rest of the sentence remains unchanged.
Write down the altered sentence and the replaced word as the output.
Output:
Meaningfulness and Correctness Prompts
Is this sentence meaningful and grammatically correct? “[Perturbed prompt]”
Answer only with Yes or No.
Table 18: Prompts used for creating perturbations. This task was preceded by a few-shot example set to guide the model in generating contextually relevant synonyms.
Classification-Based Tasks
Task Prompt template
GSM8k (0-CoT, 0-CoTB, 0-CoTR, 0-IRR, 0-PS) [Question]. [Instructional Prompt]
Write down your final answer to the question in this format: “The final answer is X.”
The type of X should be a number.
AQUA (0-CoT, 0-CoTB, 0-CoTR, 0-IRR, 0-PS) Multiple-Choice Question: [Question]
Answer Choices: [Options]
[Instructional Prompt]
Write down your final answer in the format: “The correct answer is [X].” where X is the letter of the correct answer choice (A, B, C, D, or E).
Big Bench (0-CoT, 0-CoTB, 0-CoTR, 0-IRR, 0-PS) Multiple-Choice Question: [Question]
Answer Choices: [Options]
[Instructional Prompt]
Write down your final answer in the format: “The correct answer is (X).” where X is the letter of the correct answer choice (A, B, C, D, or E).
Translation-Based Tasks
Task Prompt template
Translation (0-DSP) [Instructional Prompt]. [Original Sentence]
Before you write down the final English translation, please use these exact words: “####The final English translation of the complete sentence is:”
Translation (0-DTG) Step 1: Given the sentence: [Original Sentence], what is the English translation?
Before you write down the final English translation, please use these exact words: “####The final English translation of the complete sentence is:”
Step 2: Given the sentence: [Original Sentence], the English translation is [LLM Translation from Step 1].
[Instructional Prompt]. Before you write down the final English translation, please use these exact words: “####The final English translation of the complete sentence is:”
Table 19: Prompt templates used for both classification and translation tasks across various datasets.
A.7 Perturbations
Zero-shot
Instructional Prompt
Generated Candidates Semantic Similarity >30%
Meaningful and
grammatically correct
Final Perturbations
Let’s think step-by-step.
Let’s think slowly.
Let’s think. (removal)
Let’s think bit-by-bit.
Let’s think piecemeal.
Let’s think slowly.
Let’s think.
Let’s think bit-by-bit.
Let’s think piecemeal.
Let’s think.
Let’s think bit-by-bit.
Let’s think piecemeal.
Let’s think.
Let’s think bit-by-bit.
Take a deep breath and
work on this problem
step-by-step.
Take a deep breath and focus
on this problem step-by-step.
Take a deep breath and
on this problem step-by-step.
Take a deep breath and reflect
on this problem step-by-step.
Take a deep breath and study
on this problem step-by-step.
Take a deep breath and focus
on this problem step-by-step.
Take a deep breath and
on this problem step-by-step.
Take a deep breath and reflect
on this problem step-by-step.
Take a deep breath and study
on this problem step-by-step.
Take a deep breath and focus
on this problem step-by-step.
Take a deep breath and
on this problem step-by-step.
Take a deep breath and study
on this problem step-by-step.
Take a deep breath and focus
on this problem step-by-step.
Take a deep breath and study
on this problem step-by-step.
Feel free to ignore irrelevant
information in the problem
description.
Feel free to ignore insignificant
information in the problem
description.
Feel free to ignore
information in the problem
description. (removal)
Feel free to ignore irrelative
information in the problem
description.
Feel free to ignore unimportant
information in the problem
description.
Feel free to ignore insignificant
information in the problem
description.
Feel free to ignore
information in the problem
description.
Feel free to ignore irrelative
information in the problem
description.
Feel free to ignore unimportant
information in the problem
description.
Feel free to ignore insignificant
information in the problem
description.
Feel free to ignore information
in the problem description.
Feel free to ignore unimportant
information in the problem
description.
Feel free to ignore insignificant
information in the problem
description.
Feel free to ignore information
in the problem description.
Feel free to ignore unimportant
information in the problem
description.
Please detect the error type
firstly and refine the
translation then.
Please observe the error type
firstly and refine the
translation then.
Please the error type
firstly and refine the
translation then. (removal)
Please notice the error type
firstly and refine the
translation then.
Please discern the error type
firstly and refine
the translation then.
Please pick out the error type
firstly and refine
the translation then.
Please observe the error type
firstly and refine
the translation then.
Please the error type
firstly and refine the
translation then.
Please notice the error type
firstly and refine the
translation then.
Please discern the error type
firstly and refine the
translation then.
Please pick out the error type
firstly and refine the
translation then.
Please observe the error type
firstly and refine
the translation then.
Please the error type
firstly and refine the
translation then.
Please notice the error type
firstly and refine the
translation then.
Please discern the error type
firstly and refine the
translation then.
Please observe the error type
firstly and refine
the translation then.
Please notice the error type
firstly and refine the
translation then.
Please discern the error type
firstly and refine the
translation then.
Table 20: Illustration of the multi-stage filtering of prompt perturbations for classification prompts. Each candidate is generated via synonym, co-hyponym, or removal, and filtered for semantic similarity (>>30%) and grammaticality. Underlined candidates indicate perturbations that were rejected at a given filtering stage. The final column shows the valid perturbations used for evaluation.
Original Word Generated Candidates 20% 30% 40% 50%
Let’s We should (49.13%) Accept Accept Accept Reject
It is recommended that we (21.29%) Accept Reject Reject Reject
We can (51.57%) Accept Accept Accept Accept
First Before anything else (47.05%) Accept Accept Accept Reject
Right off the bat (25.66%) Accept Reject Reject Reject
Understand Perceive (58.85%) Accept Accept Accept Accept
Apprehend (42.68%) Accept Accept Accept Reject
Problem Dilemma (72.35%) Accept Accept Accept Accept
Hurdle (35.48%) Accept Accept Reject Reject
Difficulty (60.58%) Accept Accept Accept Accept
Devise Design (37.65%) Accept Accept Reject Reject
Draft (36.57%) Accept Accept Reject Reject
Set up (36.85%) Accept Accept Reject Reject
Plan Procedure (44.23%) Accept Accept Accept Reject
Strategy (57.06%) Accept Accept Accept Accept
Solve Crack (29.82%) Accept Reject Reject Reject
Tackle (40.75%) Accept Accept Accept Reject
Step-by-step Progressively (38.88%) Accept Accept Reject Reject
Phase by phase (30.38%) Accept Accept Reject Reject
Inch by inch (21.12%) Accept Reject Reject Reject
Accuracy 85% 95% 65% 40%
Table 21: Evaluation of semantic similarity thresholds (20%–50%) on the 0-PS prompt using 20 manually validated word variants. Each row lists an original word, its generated candidate replacements (with similarity scores), and whether the candidate is accepted or rejected at each threshold. Bold and underlined entries indicate incorrect acceptance/rejection (contradict human judgment). The accuracy row shows the overall agreement rate with human judgments for each threshold. The 30% threshold yielded the best performance (95% accuracy), which is why it was chosen for the main experiments.
A.8 Token Usage per Instructional Prompt
AQUA-RAT Big Bench GSM8K
Input tokens Output tokens Total tokens Input tokens Output tokens Total tokens Input tokens Output tokens Total tokens
0-CoT 600765 1825656 2426421 871710 1204340 2076050 488985 1098887 1587872
0-CoTB 1735827 3800599 5536426 2464368 2502721 4967089 1435263 3207869 4643132
0-CoTR 2472444 4488703 6961147 3447846 2974113 6421959 2070036 4367080 6437116
0-IRR 1390389 3014827 4405216 1974426 1831424 3805850 1149441 1895271 3044712
0-PS 4088076 9379794 13467870 5581284 6834640 12415924 3472044 7837170 11309214
Total 10287501 22509579 32797080 14339634 15347238 29686872 8615769 18406277 27022046
Table 22: Estimated token usage per instructional prompt across datasets using GPT-4o mini. Input, output, and total tokens are approximated using space-based tokenization and may differ from GPT-4o mini’s actual method.